### v1

# A Survey on Federated Analytics: Taxonomy, Enabling Techniques, Applications and Open Issues

Abstract
The escalating influx of data generated by networked edge devices, coupled with the growing awareness of data privacy, has promoted a transformative shift in computing paradigms from centralized data processing to privacy-preserved distributed data processing. Federated analytics (FA) is an emerging technique to support collaborative data analytics among diverse data owners without centralizing the raw data. Despite the wide applications of FA in industry and academia, a comprehensive examination of existing research efforts in FA has been notably absent. This survey aims to bridge this gap by first providing an overview of FA, elucidating key concepts, and discussing its relationship with similar concepts. We then conduct a thorough examination of FA, including its taxonomy, key challenges, and enabling techniques. Diverse FA applications, including statistical metrics, set computation, frequency-related applications, database query operations, model-based applications, FL-assisting FA tasks, and other wireless network applications are then carefully reviewed. We complete the survey with several open research issues and future directions. This survey intends to provide a holistic understanding of the emerging FA techniques and foster the continued evolution of privacy-preserving distributed data processing in the emerging networked society.
Index Terms:
Federated analytics, privacy, data science, security, distributed systems, federated learning, Internet-of-Things
I Introduction
There’s been a surge in data volume generated recently by the exponential growth of the Internet of Things (IoT) devices, where the number of connected devices is forecasted to reach 125 billion by 2030[1]. This IoT expansion is expected to generate 79.4 ZB of data by 2025[2]. Uploading and processing these data centrally present huge challenges for data communication and computing. Coincident with the extraordinary increase in the volume of data has also been a growing appreciation of the importance of data privacy, reflected by the enactment of stringent regulations such as the General Data Protection Regulation (GDPR) [3]. Fueled by the data explosion and privacy concerns, there is a noticeable shift from centralized to distributed data collection, storage, and processing. This evolution in data management underscores a new paradigm called federated computing, where data-oriented tasks are conducted among distributed data owners without uploading the local raw data. Federated learning (FL) is a instance of this computing paradigm that targets at collaborative model training without centralizing the raw data. It has been widely studied, deployed in real world, and adequately surveyed[4][5][6].
Federated analytics (FA) is another emerging instance of the federated computation paradigm, which conducts data analytics tasks on distributed data held by different data owners (clients). The term “federated analytics” was first introduced by Google in 2020 [7] to describe “the practice of applying data science methods to the analysis of raw data that is stored locally on users’ devices”. The first application of FA is conducted by Google researchers [7], where the FA scheme is designed to evaluate the accuracy of a trained FL model against client data in local devices. FA extends the application of currently successful federated learning (FL) to broader data analytics tasks other than model training for deep learning. In contrast to classical data analytics workflow, which usually requires the clients to upload their raw data to a centralized server, FA prevents any transmission of raw data leaving the client they originated, in order to save communication cost and preserve the data privacy of clients.
In an FA algorithm, the clients utilize their local data and the computation model received from the server to conduct local computation. The output of the local computation procedure is called “insight”. The insights are tailored data structures or information that reflect useful knowledge about the local data for the host data analytics task while preserving the privacy of the raw data. The clients upload the insight to the server to circumvent the raw data transmission, and the server performs insight aggregation to transform the individually-generated insights received from the clients to derive the population-level data analytics results. With FA, valuable analytics results can be derived to serve the pervasive data analytics needs of networked applications. The growing data privacy concerns of the users can also be properly handled.
Specifically, compared with traditional centralized data analytics and its sibling federated learning technique, FA offers the following benefits:
  • Communication reduction: FA transmits the insight of client data instead of the raw data. The structure of insight is usually designed to contain only valuable information to resolve the data analytics tasks, and the size of insight is usually independent of the number of data samples held by the clients. Therefore, FA can reduce the communication cost in data analytics, especially when the clients hold large-volume raw data with many data samples.
  • Privacy preservation: The tenet of FA prevents the transmission of raw data. The transmitted insight, which only includes abstractive information of the raw data, is less sensitive than the raw data transmission. Therefore, compared to traditional data analytics that require raw data transmission, FA enhances privacy preservation in utilizing edge data, with both the prevention of raw data transmission and additional privatization mechanisms. FA can be applied to both large-scale mobile devices, as well as several major data silos.
  • Expansive task coverage: Different from FL which mostly focuses on neural network-based predictive tasks, FA extends the scope and covers the whole spectrum of descriptive tasks in data science. These tasks range from basic calculations, such as computing averages, to more complex operations in fields like graph analytics or video analytics. FA, joining forces with FL, successfully addresses the entirety of the learning and analytics problem space, enabling all data-oriented tasks to be completed in a federated environment.
With these unique advantages, FA has offered a range of applications across various industries and sectors. For example, Google has utilized FA to improve federated learning models for its Gboard application [7] that it collects data from users’ keyboard activities to test the accuracy of its next-word prediction model. FA can also assist FL in terms of client selection [8], client clustering [9], and improving the privacy of aggregation [10]. Furthermore, sectors where data sensitivity is paramount can benefit from privacy-preserving FA to cooperate on tasks and research without sharing raw data such as healthcare[11], finance, and public services[12]. In these contexts, even competitive parties can conduct federated analytics collaboratively for common interests. The reductions in communication and computation further enable FA to offer wide applications in sectors where data is collected from remote locations like agriculture and environmental monitoring or where large-scale data collections and processing are required such as tasks in smart cities and in industry 4.0. The wide applications and studies of FA make it the proper time to survey this prominent area.
I-A Comparision and our contributions
Although the extraordinary advantage of the FA framework in privacy preservation and high-utility data analytics has spurred numerous FA studies, a systematic review of the current status of FA has not been conducted thus far. To date, only two preliminary surveys on FA exist[13, 14]. Survey [13] first reviews the federated analytics, and clarifies its position in the research literature, motivation, and application of it. However, it lacks detailed elaborations on tasks that can be supported by FA, presenting only two cases as examples. They place more emphasis on discussing the challenges, existing and possible solutions of FA in terms of the architecture, privacy problem, computation and communication resource, analytics design, and business models. Authors in [14] focus on common queries of interest in FA and their corresponding existing solutions, algorithms, and applications. They summarize and divide queries in FA into three categories: statistical queries, set-based queries, and matrix transformation queries. However, this classification is not comprehensive enough to cover other query types and data analytics tasks in FA, such as database operations, model-based queries, and tasks for FL assistance. Additionally, they do not provide a comprehensive summary of common techniques employed in FA, such as privatization, analytics, and system optimization. To the best of our knowledge, there is still no existing work to provide a comprehensive and dedicated review of federated analytics.
This survey is motivated by this gap to provide a more comprehensive summarization of FA algorithms and systems. We investigate wider classes of data analytics tasks FA applies than previous surveys. We clarify the ambiguity and discuss the overlap between FA and its related fields, showing how research in related fields exactly utilizes the idea of FA to resolve data analytics tasks with privacy preservation, and include these massive works in this survey. This survey is expected to let the readers have a comprehensive view of what the data analytics problem FA studies focus on, what are the enabling technologies of FA, and what are the opening issues of FA that require further research effort. In addition, this survey is expected to inspire researchers in related fields about the relationship and application of FA in these fields. It provides a demonstration of how FA has been applied in these fields, and how FA can be potentially utilized in these fields.
To this end, the key contributions of this article are highlighted as follows:
  • We present the first state-of-the-art comprehensive survey on FA, covering the taxonomy, key challenges, and techniques, as well as a wide spectrum of FA applications.
  • We derive a taxonomy for FA from different dimensions, including data analytics tasks, scale of clients, number of iterations, coordination model, as well as threat models.
  • We summarize enabling techniques in FA for privacy preservation, data analytics, and system optimization.
  • We summarize a wide spectrum of FA applications, including statistical metrics, set computation, frequency-related applications, database operations, model-based applications, other downstream and upstream tasks in a machine learning pipeline, and wireless network applications.
  • We identify several key open research issues and challenges in FA, covering the application scenarios, algorithm design, system optimization, and cross-layer enhancement.
I-B Structure of the Survey
Figure 1: Overview of the survey.
The structure of the survey is organized as follows. Section II provides an overview of FA, including its general workflow, a general mathematical problem formulation, and its comparison with related topics. Section III introduces our taxonomy of FA algorithms and systems. Our taxonomy covers five important dimensions that have significant impacts on its real-world applications and deployments. In Section IV, key challenges in designing and deploying FA are discussed. Section V, we introduce key enabling techniques applied in FA, including privatization techniques to realize privacy preservation, analytics techniques to derive data analytics results, and deployment/optimization techniques to enable real-world applications. In the following five sections, we introduce the vast existing FA solutions in detail, where each section covers FA solutions on a particular class of data analytics tasks: Section VI covers the computation of simple statistical metrics; Section VII covers the tasks related to frequency estimation; Section VIII covers database operations (data analytics tasks formulated as SQL queries); Section IX covers the diverse model-based data analytics problems; Section X covers data analytics tasks that are tailor-designed to assist FL systems; Section XI covers the applications of FA for wireless networks. Section XII discusses open issues and future research directions in FA, and Section XIII concludes the survey. An overview of the survey is provided in Fig. 1.
Figure 2: Architecture of FA.
II Overview of Federated Analytics
In this section, we provide an overview of the emerging FA. In Section II-A, we describe the general workflow of FA. In Section II-B, we propose a general mathematical problem formulation of FA. In Section II-C, we introduce the relationship between FA and related fields.
II-A General workflow
FA aims at conducting data analytics tasks based on the federated data held by multiple clients with privacy preservation. The fundamental principle of FA to realize privacy preservation is that the raw data held by the clients are not transmitted and exposed. In practice, the prevention of raw data exposure is not sufficient to realize a formal privacy guarantee, and extra privatization mechanisms are usually applied in FA for the formal privacy guarantee. As is demonstrated in Fig. 2, the typical architecture of FA (server-client model) includes a server, which is hosting a data analytics tasks, and multiple clients, which would like to contribute their personal data to the data analytics task but require privacy preservation on their data. The right part of Fig. 2 demonstrates the steps of one-shot and iterative FA systems. In the FA workflow, the clients are in charge of a part of the data analytics task. They receive the computation model from the server, and perform local computation based on the received computation model and the client’s local data. The result of local computation, which is termed as “insight” in this survey, is uploaded to the server, where the insights are processed by privatization mechanisms to realize privacy preservation. The server aggregates the received insights with a dedicated aggregation algorithm. The aggregation output may be directly the results of the data analytics task when the FA algorithm is one-shot. When the algorithm is iterative, the aggregation result updates the computation model, and the data analytics results are derived after many rounds of model distribution, local computation, insight upload, and insight aggregation.
II-B Mathematical definition
FA considers a system of $n$ data owners, or clients, denoted $DO$, and each client $DO_{i}$ holds local client data $x_{i}$. These clients would like to collaboratively accomplish a data analytics task, which can be reformulated as deriving the value of the task function
$\mathcal{F}(x_{1},...,x_{n}).$ (1)
During the procedure of deriving $\mathcal{F}$, the personal raw data $x_{i}$ should not be exposed to any entity other than $DO_{i}$.
The aforementioned problem is resolved by the FA paradigm by designing the appropriate insight derivation function $\mathcal{I}_{i}$ and a global aggregation function $\mathcal{A}$. The clients derive insights from their local data using the insight derivation function $\mathcal{I}_{i}$, and then aggregate the insights to derive the value of the task function, i.e.,
$\mathcal{F}(x_{1},...,x_{n})\leftarrow\mathcal{A}(\{\mathcal{I}_{i}(x_{i});i\in DO\}).$ (2)
In the majority of existing FA studies, the FA algorithm treats all the data owners equally, and all the clients share the identical insight derivation function (denoted $\mathcal{I}$), and the FA procedure can be written as
$\mathcal{F}(x_{1},...,x_{n})\leftarrow\mathcal{A}(\{\mathcal{I}(x_{i});i\in DO\}).$ (3)
Based on the aforementioned mathematical formulation, the research problem in designing the FA algorithm is derived: the researchers need to design insight derivation function $\mathcal{I}$ and aggregation function $\mathcal{A}$, so that the FA result $\mathcal{A}(\{\mathcal{I}(x_{i});i\in DO\})$ should be identical (or close to) the result of the task function $\mathcal{F}(x_{1},...,x_{n})$.
The previous formulation describes the one-shot FA setting. In the iterative FA setting, the insight derivation function should take the computation model as input, and the results of the aggregation function should be the data analytics result in the last iteration and should be updated computation model in other iterations.
II-C Comparison with related topics
TABLE I: Difference between FL and FA
FL FA
Goal Training neural networks Non-training tasks
Aggregation FedAvg Task dependent
Insight Model weights Task dependent
TABLE II: Difference between DDM and FA
DDM FA
Raw data redistribution & transmission Allowed Forbidden
Clients and server Trusted Untrusted
Heterogeneities Little concerned Focused
Although the term FA has been introduced in recent years, some core characteristics of FA, such as privacy protection, distributed computing, can already be found in existing research fields, such as FL, privacy-preserving data mining, and distributed data analytics. These topics naturally have some overlap with FA. We list these three related topics of FA, introduce the overlapping research, and discuss how FA extends their concepts.
II-C1 Relationship between FL and FA
Both Fl and FA target at conducting intelligent data-oriented tasks without centralizing the raw data. Their major difference can be summarized in Table. II. First, FA and FL are differentiated for their goal tasks. In FL, the mechanism is designed to serve a predictive model training task. In most of the FL literature, the “model training” in FL exactly refers to deep supervised learning on a neural network, if no specific context is present. On the contrary, as its name refers, FA tackles the extensive field of descriptive data analysis tasks, ranging from the simplest task of calculating the average to complex tasks such as graph analytics and video analytics. Second, the task range of FA diversify the forms of local insights and central aggregation. In contrast to the model weights or gradient distilled from the local clients, the insights from FA could be much more diverse and task dependent, such as privatized hashing results [15], one-bit response to server query [16], or local clustering centroids [17]. The procedure of insight aggregation of FA is also diverse and task dependent compared to FL, such as tree aggregation [18], Bayesian-based distribution estimation [19], or secret sharing decryption [20].
II-C2 Relationship between FA and privacy-preserving data mining
Privacy-preserving data mining is a comprehensive term including all data analytics schemes that take privacy concerns into consideration. Therefore, FA can be considered as a subfield of privacy-preserving data mining. As the need to perform data analytics with privacy preservation naturally exists, many solutions for privacy-preserving data analytics existed even earlier than the term federated computation (FL was first proposed in 2016 and FA was in 2020). These works (represented by RAPPOR [15]) do not use the term “federated”, but follow the similar idea and methodology as FA. Therefore, it is reasonable and scientific to re-classify these works as FA. On the other hand, not all privacy-preserving data mining research can be considered as FA. FA centers around distributed data setting and local privacy preservation. Heterogeneity and the resulting issues born at the client side, such as client dynamics, adversarial attacks, communication overhead, and incentive issues, are all new problems that are not considered in classical privacy-preserving data mining.
II-C3 Relationship between FA and distributed data mining
Distributed data mining (DDM) refers to the procedure of conducting data analytics tasks while the data is distributed among multiple parties or machines. According to such a comprehensive definition, FA is a subfield of DDM because of its federated data setting. However, the FA studies differ from the general distributed data mining in that distributed data mining usually operates in trusted database scenarios with distributed machines supporting computing acceleration. In summary, in DDM, the computation nodes are naturally trusted, and the raw data transmission is allowed, which is not allowed in FA. Table II concludes the differences between FA and DDM.
TABLE III: Examples of FA algorithms and systems following the taxonomy of this survey.
Reference Data analytics task Client scale Num. Iterations Coord. model Threat model
[21, 22, 23, 24] Mean computation $>$1,000 One-shot Server-client Model of LDP
[25, 26, 27, 28, 29] Mean computation $>$1,000 One-shot Server-client Model of cryptography
[30] Median/percentile computation 2 Iterative Decentralized Model of cryptography
[31] Median/percentile computation 200 One-shot Server-client Model of cryptography
[32] Median/percentile computation 15 Iterative Decentralized Hiding raw data
[17] Clustering 200 One-shot Server-client Hiding raw data
[33, 34, 35, 36] Clustering 4$\sim$100 Iterative Server-client Hiding raw data
[15, 37, 38, 39, 40] Frequency oracle/Frequent pattern mining $>$1,000 One-shot Server-client Model of LDP
[41, 42, 16] Frequency oracle/Frequent pattern mining $>$1,000 Iterative Server-client Model of LDP
[18, 43] Frequency oracle/Frequent pattern mining $>$1,000 Iterative Server-client Model of CDP
[44] Frequency oracle/Frequent pattern mining $>$1,000 One-shot Server-client Models of cryptography & CDP
[45] Frequency oracle/Frequent pattern mining $>$1,000 Iterative Server-client Models of cryptography & CDP
[46, 47, 48] Query 2$\sim$3 One-shot Decentralized Model of cryptography
[49] Query 2 One-shot Decentralized Models of cryptography & CDP
[50, 51, 52] Query 10$\sim$16 One-shot Server-client Model of cryptography
[53, 54, 55] Query $>$1000 Iterative Server-client Model of cryptography & CDP
[56] Matrix factorization 610 One-shot Server-client Model of cryptography
[57] Matrix factorization 40 Iterative Server-client Model of cryptography
[58] Matrix factorization 2 One-shot Decentralized Model of cryptography
[59, 60] Multi-armed bandit 10 Iterative Server-client Hiding raw data
[61, 62, 63] Multi-armed bandit 10 $\sim$100 Iterative Decentralized Model of CDP
III Taxonomy of Federated Analytics
Given the diversity of the existing FA algorithms and systems, it becomes beneficial to provide methods to classify FA problems and solutions. In this section, we provide the taxonomy of FA. It examines the design and application of FA from five important dimensions, including data analytics tasks, scale of clients, number of iterations, coordination models, and threat models.
III-A Data analytics tasks
As the application of FA covers massive and diverse data analytics needs, classifying FA algorithms and systems based on the types of data analytics tasks they solve becomes a natural and useful taxonomy of FA. The host data analytics task heavily influences the algorithmic design, characteristics, and challenges of the corresponding FA algorithms and systems, and the algorithmic design of FA is expected to be quite different when handling different types of data analytics tasks. One of its reasons is that different data analytics tasks extract and analyze different features form the raw data, which then determine the structure of required insight, the insight generation approach, and insight aggregation approach. Some researcher would like to break the limit of data analytics task-specific FA design, and propose some unified FA framework. However, proposing such an FA framework that can handle all data analytics tasks is still far from reality at this moment. Therefore, the host data analytics task is currently an important component of any FA taxonomy. In Section XII-C, we investigate the probability of a unified FA framework as a future research direction.
In this survey, we follow the taxonomy of data analytics tasks to introduce the existing FA research. We discuss five classes of data analytics applications, in Sections VI-X respectively, and introduce the existing FA algorithms and systems in the context of the host data analytics tasks.
III-B Scale of clients
FA algorithms and systems have diverse settings regarding the scale of participating clients, from fewer than 10 to more than a million. The difference has a large relationship with the privacy mechanisms they use as well as their application scenarios. For mobile applications, an FA scheme is usually expected to maximize its capacity to support as many participating clients as possible. For industrial scenarios with only a few data silos, an FA scheme is also expected to work, usually with higher requirements on analytics performance.
Usually, DP-based mechanisms are better at supporting massive participating clients, because they only require the clients to individually perturb their client uploads in parallel, and the server aggregation function usually has linear complexity [21, 22, 15, 37, 16]. In addition, when the privacy model of CDP is considered, increasing the number of participating clients can increase privacy preservation by better hiding the information of individual clients [18, 43].
On the other hand, FA schemes utilizing computation-intensive cryptography tools, though can provide higher analytics utility, are usually hindered by the algorithm, computation, or communication limits. Consequently, they can only be applied to a smaller scale of clients. For example, the garbled circuit [64], a classical and famous cryptography tool employed in some FA studies [46, 50], only supports computation between two parties, and modifications of it usually introduces significant computation and communication overhead to support several more participating clients. To overcome the limit in supported client scale when utilizing cryptography tools, many research efforts are proposed to increase the supported participating clients in cryptography-based FA systems. An example direction of these research is to replace the computation-intensive cryptography tools by lightweight ones, like simple additive masking [25, 26, 27, 28].
III-C Number of iterations
FA algorithms can be classified into one-shot (one-round) non-iterative schemes and iterative schemes. The one-shot schemes are completed after conducting model distribution, local computation, and insight aggregation once. In contrast, the iterative schemes repeat the phases, constantly deriving new knowledge and optimizing the computation model, until the data analytics task is accomplished. Compared to the one-shot algorithms, iterative algorithms usually require more complex design, because iterative algorithms requires extra considerations about how to maintain the “state” of the data analytics procedure, and how the local computation procedure of the clients can be guided by the global knowledge derived in previous rounds. Roughly speaking, one-shot schemes are usually applicable in simple FA tasks and incur less communication cost, like federated mean computation [21, 22], while iterative schemes are usually required for complex FA tasks with explicit computation models, like federated multi-armed bandits [59, 60, 61, 62, 63].
III-D Coordination models
The vanilla setting of FA considers the server-client architecture, where all the clients only communicate with the centralized server. The server-client architecture becomes the mainstream form of FA because the clients can easily register themselves in the system, the FA system can easily scale up, and the communication scheme can be easily determined. However, as some shortcomings of the server-client architecture are observed by the researchers, alternative architectures, such as server-client architecture with client-client communication [25, 26], and even a fully decentralized architecture without any centralized server [61, 62, 63], are proposed. These alternatives have potential advantages such as preventing one-point failure, better privacy preservation, better communication efficiency, and adaptive organization.
III-E Threat models
While FA algorithms and systems apply different privatization techniques to guarantee privacy preservation (Section V-A provides a demonstration), different threat models are considered that correspond to these privatization techniques. For example, in these FA systems utilizing cryptography tools like homomorphic encryption or multi-party computation, the threat model can be assumed that privacy is preserved when the server cannot know the decrypted upload from the clients; in these LDP-based FA schemes, the threat model prevents the adversary from inferring the client raw data by analyzing the client upload; in these CDP-based FA schemes, the threat model prevents the adversary from inferring the client raw data by analyzing the final analytics results. There does not exist an inclusion relationship between the threat model of cryptography tools and the threat model of DP, and one cannot claim that one is strictly stronger than the other. In addition, FA threat models can also be classified regarding the trust in the server or aggregator. Few works assume that there exists a trusted server or aggregator; the majority of FA works assume that the server is honest-but-curious (also known as semi-honest); the rest of FA research investigates privacy reservation over stronger adversarial models.
The previous taxonomy with five dimensions provides us with a powerful tool for grouping and classifying FA algorithms and systems. In Table III, we list some existing FA solutions, and show their positions according to our taxonomy. [^1] Throughout the survey, the client scale, or maximal supported number of clients, is counted as the maximal number of clients in these papers’ experiment settings.
IV Key challenges in Federated Analytics
As FA is expected to handle various data analytics tasks over federated data with strong privacy preservation, there exist many challenges for researchers to realize the ambition of FA. In this section, we investigate some dimensions of FA challenges and discuss how these challenges could be handled by novel FA research attempts.
IV-A Accuracy, privacy, and performance trade-off
FA aims to extract and analyze insights from decentralized data sources in a privacy-preserving way. There are trade-offs as regards accuracy, privacy, and performance, which are three of the most important aspects of FA. Researchers in [49] pointed out these trade-offs among differential privacy, secure multi-party computation, and approximate query processing. Firstly, when discussing trade-offs between accuracy and privacy, it is imperative to bring up the differential privacy (DP) which has been a widely acknowledged privacy-preserving criterion in data analytics. The details of DP are described in Section V-A herein. To satisfy DP, mechanisms usually introduce random noise into a result. An increase in noise addition corresponds to heightened privacy protection since it more and more obfuscates the contribution of each individual. However, the guarantee of privacy comes at the cost of a reduction in accuracy, and the result deviates from the ground truth due to the injection of noise. Secondly, the trade-off including performance is usually related to another sort of privatization technique related to cryptography. Although it guarantees both privacy and accuracy by encrypting the data of each individual before sharing it and obtaining the real result by corresponding decryption, there is an enormous amount of computation and communication overhead when performing such operations in FA. Lastly, there are approximating methods applied in FA to increase the performance of the system which sacrifices the accuracy of results. For example, private sampling algorithms utilize less data in the computation and communication process while introducing variance from sampling into the final result. In conclusion, balancing these trade-offs is a key challenge in FA.
IV-B Intensive computation
In FA, cryptography is commonly used to protect the privacy of each client’s data while it is processed and analyzed by multiple parties including the aggregator and other clients. However, cryptographic methods are considerably resource-intensive, both in terms of computation and communication. Operations in cryptography like encryption, decryption, and hashing are computationally intensive. For some advanced cryptographic techniques used in FA like homomorphic encryption and secure multi-party computation which allow various computations on the encrypted data, the required computation resource is even higher [65]. Regarding communication cost, encrypted data tends to be larger than their plaintext counterparts. Furthermore, most cryptographic protocols require multiple rounds of communication between parties [10], exchanging messages like intermediate results. This attribute of cryptography extremely increases the bandwidth required for tasks in FA. In addition, managing cryptographic keys is also a complex task in federated settings. They are required to be private and secure in their generation, distribution, rotation, and destruction [66]. This process will be more complex and computationally intensive if the number of clients increases [29]. These attributes together contribute to the large resource demands for utilizing cryptography in FA, which is a challenge to preserve privacy efficiently in terms of computation costs.
IV-C Wireless communication
Wireless communication becomes a major challenge in FA primarily due to its distributed and dynamic nature. The slowest participant in this networked system can significantly hinder overall performance, known as the “straggler effect”. This is often exacerbated by varying data transmission rates and unreliable communication channels, leading to delays and inconsistencies in data aggregation. The paper [67] addresses it in FL by clustering user equipment based on upload times, thereby reducing time divergence among participants. In addition, the over-the-air (OTA) [68] technique brings efficiency as well as new challenges to federated analytics and federated learning, such as signal interference and the requirement for precise synchronization. Furthermore, the mobility of devices in wireless networks, such as in vehicular networks and aerial networks, introduces another layer of complexity, requiring adaptive strategies for data transmission and model training that can cope with changing network typologies and varying channel conditions. The integration of technologies like Reconfigurable Intelligent Surfaces (RIS) shows promise in enhancing communication efficiency and reliability in these scenarios [69], while they also require sophisticated optimization of resources and careful handling of estimated Channel State Information (CSI).
IV-D Limited privacy resource
Privacy resource in the context of FA usually refers to the “privacy budget” in the DP mechanisms, which intuitively means the amount of allowable privacy loss for each individual within an analysis of the whole data set. Although the information of clients is protected by various privacy-preserving techniques, there is still a risk that repeated analyses such as multiple or iterative queries can expose it. Typically, a portion of the privacy budget is consumed and a specific amount of noise will be added to the results for a query. Managing the privacy budget is important since multiple analyses may be performed on the same group of clients. Every analysis incrementally raises the risk of privacy loss so this cumulative effect ought to be monitored and controlled. How to effectively schedule the privacy budget for the multi-query analysis is a challenge in FA concerning the accuracy of results for each query and total privacy loss for each client.
IV-E Robustness
The robustness challenge in FA encompasses a range of considerations that are similar in FL, including malicious and non-malicious data corruption. Malicious data corruption involving deliberate attacks by malicious parties is a considerable threat to the integrity and security of FA systems. These attacks such as data poisoning and privacy breaches call for corresponding robust security measures, such as encryption and authentication to discover and defend. On the other hand, non-malicious data corruption caused by unintentional errors and client dropout due to either connection issues or device failure introduces bias to influence the accuracy of analytics or even causes failure of the system. In this context, there’s a requirement for fault-tolerant mechanisms to cope with client dropouts for robustness, ensuring that the system works effectively even with such disruptions. Addressing these challenges is crucial to maintaining the robustness of FA systems in terms of reliability, security, and accuracy.
IV-F Data heterogenity
Data heterogeneity, i.e., the phenomenon that data possessed by different clients follows heterogeneous distributions, is common in federated systems. As the tenet of privacy preservation restricts raw data from being transmitted, data reallocation or adjustment is usually unavailable for federated systems, making data heterogeneity unavoidable. The challenge of data heterogeneity has become one of the most popular research focuses in the field of FL [70], and also widely exist in FA systems. Since FA systems tackle various data analytics tasks, data heterogeneity introduces diverse influences for different heterogeneous tasks. In the federated MAB model [59, 60, 62], data heterogeneity is formulated as the biased reward of the same arm from different clients, and FA schemes are proposed to guarantee the performance of the model globally or personally. In [17], data heterogeneity is studied in a federated clustering problem, where the data points held by one client are biased to one or several clusters. The authors of [17] found that the data heterogeneity is beneficial for the FA scheme, where a higher analytics performance can be achieved in the heterogeneous data environment. In [8], an FA scheme is proposed to measure the severity of data heterogeneity directly. The proposed scheme can figure out clients with low/high data heterogeneity, which provides significant information to other federated tasks in handling the data heterogeneity challenges.
IV-G Fairness
The challenge of fairness has been investigated in many federated systems, mostly for FL. As the term “fairness” is a comprehensive term. Researchers investigate multiple interpretations of fairness, based on different types of benefits focused by the players. The existing works mainly consider three dimensions of fairness. 1) Collaboration fairness, where some monetary rewards or their equivalents are assumed to be awarded to the participating clients after the federated model is trained, and the participating clients would like to fairly allocate such benefits based on their contribution to the federated system [71, 72, 73, 74]; 2) Model performance fairness, where the finally derived federated model is adjusted to obtain a fair and balanced performance among the clients with heterogeneous data [75, 76, 77]; 3) Participation fairness, which focuses on balancing the chance of clients to become participants of the federated system [78].
Although existing fair federated system research mainly focuses on FL, the rationale and fairness concepts are still applicable to FA algorithms. In addition, the novel characteristics of FA, including the vast investigated data analytics problems, may raise new fairness concepts and challenges, which require tailored solution concepts.
V Key Techniques in Federated Analytics
In this section, we investigate the key enabling techniques that are frequently utilized by FA. These key techniques are categorized by their utilization in FA: In Section V-A, the privatization techniques, i.e., those enhance the privacy preservation of FA, are discussed; in Section V-B, the analytics techniques, i.e., those help derive the analytics results, are presented; in Section V-C, the deployment and optimization techniques, which improves the system availability, are presented
V-A Privatization techniques
The tenet of FA prevents the exposure of raw data to any entity other than the data owner (client), which naturally provides privacy preservation to some extent. However, such a “hiding raw data” idea fails to provide any privacy guarantee. Therefore, existing FA research prefers to enhance the privacy preservation of their schemes by providing some formal and rigid privacy guarantee, by leveraging various privatization techniques. In this part, we analyze the privacy guarantees and the corresponding techniques to realize these techniques.
Local differential privacy Differential privacy (DP), introduced by [79], is a privatization technique enforced on any data publication mechanism. It has become the most popular criterion for privacy preservation in data analytics, as well as FA. It considers the threat models where the data publication mechanism publishes some outputs based on the raw data, and an adversary tries to infer sensitive information based on the output, such as membership inference attacks [80] and data reconstruction attacks [81]. DP handles such threat models by introducing randomness to the output of the data publication mechanism, and restricting the probabilistic distribution of outputs.
Local differential privacy (LDP) is the most widely used DP variation in FA, as it provides the strong local privacy preservation required by the majority of FA applications. LDP-based FA schemes consider the threat model that the clients upload their insights to an untrusted server, where the untrusted server directly leverages the received insights to perform inferences. LDP is enforced by applying randomness in client uploads, so that the probability of any output would not change intensely for clients with arbitrary different local data. The definition of LDP in an FA setting is given as follows.
Definition 1 (Local differential privacy)
Consider a data publication scheme $\mathcal{M}$ executed by the clients in an FA system. Denote $\mathcal{D}$ as the set of all possible client data. $\mathcal{M}$ satisfies $(\epsilon,\delta)$-LDP when
$\mathbb{P}\big{(}\mathcal{M}(d_{1})=x\big{)}\leq e^{\epsilon}\mathbb{P}\big{(}\mathcal{M}(d_{1})=x\big{)}+\delta,$ (4)
for any $d_{1},d_{2}\in\mathcal{D}$, and any possible output $x$ of $\mathcal{M}$.
$\epsilon$-LDP, a stronger and stricter LDP criterion, can be achieved by setting $\delta=0$ in (4). In the setting of LDP, the data held by a client is considered as a data sample $d\in\mathcal{D}$, and LDP ensures that the probability of deriving any upload will not change intensely for clients possessing different data samples. An FA scheme reinforced by LDP can guarantee that the insight uploaded to the server cannot be leveraged by an adversary to efficiently infer the raw client data, from the probability theory perspective.
The satisfaction of LDP requires perturbation (randomness) on the client uploads. For the binary data, LDP is typically realized by the randomized response, i.e., flipping each bit with some probability. The randomized response-based LDP mechanisms usually work in FA applications where the local computation results are naturally in the binary form [21, 16, 38, 40] or the client data are encoded into a binary form with sketching techniques [15, 42, 82]. For the continuous scaler data, LDP is typically realized by introducing continuous scaler noise on the data. Gaussian mechanism and Laplace mechanism are the most widely used schemes to realize LDP for continuous scaler data. Gaussian mechanism is applied by adding a random noise following the Gaussian distribution into the uploaded scaler, which can guarantee the corresponding $(\epsilon,\delta)$-LDP [83, 84]. Laplace mechanism, introducing a random noise following the Laplace distribution, is able to satisfy the stronger $\epsilon$-LDP, which have been utilized in [23, 85, 55].
Central differential privacy Central differential privacy (CDP), also known as global differential privacy, is the original form of DP. It considers the scenario that a data analytics mechanism processes a database containing many data samples, and the published analytics results can protect the raw data of any single data sample from being inferred by an adversary. The definition of CDP is as follows.
Definition 2 (Central differential privacy)
Consider a data publication mechanism $\mathcal{M}$. Denote a pair of neighboring databases $\mathcal{D}$ and $\mathcal{D}^{\prime}$ that differs in at most one data sample. $\mathcal{M}$ satisfies $(\epsilon,\delta)$-CDP when
$\mathbb{P}\big{(}\mathcal{M}(\mathcal{D})=x\big{)}\leq e^{\epsilon}\mathbb{P}\big{(}\mathcal{M}(\mathcal{D}^{\prime})=x\big{)}+\delta,$ (5)
for any $D,D^{\prime}$, and any possible output $x$ of $\mathcal{M}$.
Similarly, $\epsilon$-CDP can be achieved by setting $\delta=0$ in (5). When enforcing the same privacy parameters, CDP usually requires much smaller noise than LDP, which can improve the correctness of the final analytics results.
Compared to LDP, where the client data are considered data samples and directly protected, CDP cannot be directly applied to many FA scenarios because it considers enforcing protection on the data already aggregated. In the FA literature, there are mainly two settings to enforce CDP. In the first setting, the raw data of each client is considered as a database consisting of multiple data samples. The clients enforce CDP on the database, so that information about each individual sample is protected. FA research considering this setting includes [86, 49]. In the second setting, similar to the setting of LDP, the data within a single client is considered as a data sample in $\mathcal{D}$. The CDP scheme guarantees that an adversary cannot infer the sensitive information of a client by analyzing the final FA result. However, such privacy attacks can succeed when the adversary has access to the uploads of the clients (e.g., when the server acts as the adversary). Therefore, the second setting usually assumes a trusted aggregator (server), and sacrifices the formal privacy guarantee. The second setting has been considered in [87, 88, 18, 43].
These exist diverse techniques to realize CDP. Similar to those used for LDP, the Gaussian mechanism can still be applied for CDP to protect continuous scaler output, which can satisfy $(\epsilon,\delta)$-CDP [86, 49, 87, 88]. [^2] Laplace mechanism also has theoretically applicable for CDP, but is not selected for researchers due to the high variance of Laplace noise. Compared to the LDP case, where Gaussian noise is added to each single data sample, applying the Gaussian mechanism for CDP only requires adding noise into the aggregated result. Therefore, the overall noise on the final analytics results is significantly reduced. The exponential mechanism tackles the analytics tasks where a single sample in the database serves as the output (e.g., the federated median computation task). It realizes $\epsilon$-CDP by calculating a utility score (preference of serving as the output) for each data sample, and then selecting the output data sample based on the probabilistic distribution characterized by the utility scores. The exponential mechanism has been applied to realize CDP and also distributed differential privacy, which will be introduced later. [89] Conducting sampling on the data samples in the database is another effective tool to realize CDP. When only a random portion of data samples are selected for the data analytics task, randomness is introduced to the final analytics results, which can satisfy the criterion of CDP in (5) after proper mathematical proof. Two approaches are considered to utilize the sampling technique. In [18, 43], sampling is applied to directly realize CDP without any noise injection procedure; in [49, 55], sampling can reinforce a scheme that already satisfies DP via noise injection, by deriving the stronger privacy preservation (decreasing the $\epsilon$ parameter). The latter approach can be applied for both reinforcing CDP and LDP.
Distributed differential privacy As mentioned above, LDP can achieve a stronger local privacy guarantee, without the assumption of a trusted aggregator; CDP introduces randomness with a smaller magnitude, which derives a higher correctness in final analytics results. Distributed differential privacy (DDP) is an attempt the combine the advantages of LDP and CDP. DDP statistically satisfies CDP to privatize the aggregated result. In addition, DDP requires clients to utilize cryptography tools to encrypt their uploads. With the encryption, the aforementioned local privacy risk existing in CDP, that an adversary can infer sensitive information when observing the individual client upload, is then eliminated.
The deployment of DDP consists of two phases: distributed noise generation and encrypted aggregation. In the distributed noise generation phase, each client injects noise on their local uploads. The noise is deliberately designed so that the aggregated noise of the clients can satisfy the criterion of CDP. In the encrypted aggregation phase, the client upload, possessed by distributed noise generation and uploaded to the server with encryption. The server cannot observe the individual upload of any client, but can only observe the CDP-enforced aggregated uploads.
DDP algorithm designers can freely choose and combine the techniques for distributed noise generation and encrypted aggregation. Many options for these techniques are surveyed in [90]. DDP has been applied in FA systems, like [12, 89, 44, 45].
In Fig. 3, we summarize the three kinds of DP variations utilized by FA, and the corresponding techniques to realize these DP variations.
Refer to caption[15, 21, 16, 38][42, 40, 82][83, 84][23, 85, 55][86, 49, 87, 88][89][18, 43][49, 55][12, 89, 44, 45]
Figure 3: Summery of DP variations, their enabling techniques, and existing works applying them. Sampling is not only able to solely satisfy CDP, but also can enhance the privacy preservation of any existing CDP/LDP scheme. Applications following the latter approach are marked red in the figure.
(a) Cryptography-based aggregation
(b) $k$-anonymity
(c) LDP
(d) DDP
Figure 4: Summery of some privatization techniques in FA
$k$-anonymity The concept of $k$-anonymity emerged as a privacy-preserving technique, specifically crafted to safeguard individual data in publicly accessible data sets, addressing the growing concerns over personal data security and privacy breaches [91]. The main idea is to ensure that any specific individual cannot be linked given any record in the dataset by guaranteeing that any combination of attribute values will indicate at least $k$ individuals. In other words, for every individual in the data set, there are at least $k-1$ other members with the same information thus it’s rarely possible to search out an individual based solely on certain attribute values. To achieve $k$-anonymous, some attributes are either removed or suppressed. There are three types of attributes in datasets: identifiers, non-identifiers, and quasi-identifiers. Quasi-identifiers are attributes that can potentially distinguish individuals together such as age, zip code, and gender, none of which can identify an entity alone. Generally, identifiers are suppressed and quasi-identifier are either removed or generalized such as replacing the ages with age ranges to achieve $k$-anonymous.
It’s a suitable tool in federated analytics in terms of preventing the leakage of sensitive information and the identification of individuals. Usually, $k$-anonymity is utilized in phases before sharing information. The risk of privacy breaches is reduced by guaranteeing that the data set of each client follows k-anonymity before sharing with the server or each other. In [46], they designed the Private Data Network to enable querying in databases of different parties without disclosure of their raw data by appling $k$-anonymity before data sharing. With the trusted third party as the query planner and results collector, the $k$-anonymous data of clients are then queried and aggregated to solve the questions from users. This setting is particularly relevant to collaboration in clinical research of healthcare clients. There are other methods to achieve $k$-anonymous. Researchers in [61] utilized sampling and trie pruning to fulfill $k$-anonymity for obtaining DP without noise injection. In [20], authors applied a more straightforward method that only the data sent by more than $k$ clients are chosen in the aggregation phase.
Homomorphic encryption Homomorphic encryption is a powerful privacy-preserving cryptographic method allowing to perform mathematical computations such as addition and multiplication on encrypted data without decryption. It converts original data to cyphertexts by a public key, which cannot be read before it’s transformed back with the corresponding private key. Homomorphic encryption plays an important role in federated analytics in that it enables the aggregation of encrypted data without access to clients’ plaintext data. The collected encrypted data are then performed with mathematical operations to obtain a result which is also encrypted with support of its homomorphic attribute. The outcome is usually sent to a trusted third party holding the private key for decryption and is finally revealed to the querier. In this way, clients cooperate on data analysis tasks but raw data never leaves the local position of their owners. There are several types of homomorphic encryption, including fully and partially homomorphic encryption.
Homomorphic encryption was early applied in private set intersection tasks[92, 93, 94, 95, 51]. Generally, in the two-party setting, the sender and receiver first agree on a homomorphic encryption scheme and the receiver generates a public-private key pair. The receiver then sends the encrypted set by public key to the sender who compares it with her encrypted set and sends the intersection back. Finally, the receiver revealed the encrypted intersection entities by the private key. The homomorphic attribute plays a crucial role in the faster computation of the phase in which the sender identifies common encrypted elements. Homomorphic encryption is also a suitable tool when there’s a requirement for allowed mathematical operations such as total summation in aggregation parts. In [29], researchers proposed Honeycrisp to solve the count mean sketch task at a large scale without a trusted core. A committee randomly selected from users generates keys and each client sends their additively homomorphically encrypted data by “two-element” Ring-LWE-based encryption to the server which obtains the total encrypted sum and sends it to the committee for decryption. Authors in [56] proposed FedMF, a secure matrix factorization framework that also used an additive homomorphic encryption, Paillier encryption, to secure the gradient information of each user.
TABLE IV: Summary of homomorphic encryption.
Reference Task Method Type
Freedman et al. [94] PSI Paillier encryption partially
Huang et al. [95] PSI ElGamal encryption partially
Freedman et al. [93] PSI ElGamal encryption partially
Chen et al. [92] PSI FHE with fine tuned parameters fully
Zhang et al. [51] PSI ElGamal encryption partially
Honeycrisp [29] Count mean sketch “two-element” Ring-LWE-based encryption fully
FedMF [56] Matrix factorization Paillier encryption partially
Secure multi-party computation Secure multiparty computation (MPC) aims to jointly compute functions with inputs from several parties while guaranteeing the privacy of these in- puts. MPC ensures privacy and security in federated analytics by allowing multiple parties to collaborate on tasks without compromising the confidentiality of individual data sets of each party. Secure multi-party computation focuses less on the requirement for a large volume of data to maintain privacy and accuracy when compared to differential privacy. Nevertheless, it imposes a heavier burden in terms of computational and communication costs. While it can excel in preserving privacy and ensuring precise results with any data size, the trade-off lies in the heightened demand for computational resources and communication overhead among participated parties, rendering it a more resource-intensive approach. Some cryptographic techniques, like homomorphic encryption, garbled circuits, or secret sharing, are usually utilized to enable secure computation among parties. The choice of technique depends on the specific requirements and constraints of various federated analytics tasks. In the computation phase, only intermediate results are shared instead of users’ raw data, which are also usually encrypted to guarantee the throughout the whole process.
MPC has a broad application in federated analytics. It’s used in tasks like secure communication of the median [30, 31], multi-party set intersection [96], and searching heavy hitters [44]. In [30], it focuses on securely computing the kth-ranked element (such as the median) of combined confidential datasets from multiple parties. The protocols employ techniques like binary search and consistency checks to ensure privacy while computing the median. While [31] addresses distributed private learning and focuses on computing the median in a way that is both differentially private and efficient in terms of computation and scalability. It utilizes a multi-party computation (MPC) approach to compute the exponential mechanism for the median, which is also adaptable for other rank-based statistics and machine learning optimizations. In [96], the authors introduce an efficient method for multi-party private set intersection (PSI), allowing multiple parties to compute the intersection of their datasets without revealing extra information. It uses a new paradigm based on oblivious programmable pseudorandom functions (OPPRF) and avoids computationally expensive public-key operations. Also, the protocol is secure against colluding semi-honest parties and has been demonstrated to be practical for up to 15 parties, each with datasets of a million items. Researchers in [44] present efficient multi-party computation (MPC) protocols, HH and PEM, to compute differentially private heavy hitters. These protocols use sketches for approximate counts, offering better accuracy than local differential privacy, and are practical for different data sizes: HH is suitable for small datasets (running time linear in data size), while PEM is more efficient for larger datasets (running time sublinear in data domain). The protocols are more accurate than local differential privacy approaches and an accessible performance on computation and communication cost. Also, some MPC protocols or platforms are designed to perform federated analytics [47, 97, 50, 51, 52]. Conclave [47] is a query compiler designed to accelerate relational analytics queries by combining data-parallel local cleartext processing with smaller secure multi-party computation (MPC) steps. It offers a hybrid MPC-cleartext protocol for cases where parties trust others with specific subsets of data. Cerebro [97] is an end-to-end collaborative learning platform that enables multiple parties to compute learning tasks without sharing plaintext data. It addresses the need of organizations to collaboratively use sensitive data, complying with policy regulations and business competition. Cerebro balances system design for safe collaboration with release policies and auditing, aiming to simplify the complex performance trade-offs between different MPC protocols. Authors in [50] introduce Senate, a system for MPC that allows multiple parties to collaboratively execute analytical SQL queries without exposing individual data, even with the presence of malicious adversaries. The main idea of Senate is MPC decomposition which enhances computational efficiency by breaking down cryptographic computations into smaller, parallel units. In [51], researchers focus on advancing the field of skyline queries, which are crucial for multi-criteria decision-making systems. It proposes a local dominance-based framework to enhance the efficiency of skyline queries in a vertical data federation setup. The framework decomposes skyline queries into more manageable units, improving the overall query process’s efficiency and security. Hu-Fu [52] is the first system dedicated to efficient and secure spatial query processing in a data federation context by optimizing the balance between plaintext and secure operations, minimizing the use of secure operators while still maintaining the overall security of the process. It’s a significant advancement in data federation that Hu-Fu parses federated spatial queries written in SQL, decomposes them into secure and plaintext components, and securely collects query results, offering a practical solution to the challenges of processing spatial queries securely and efficiently. Besides, authors in [29] designed a protocol to distribute the secret key of homomorphic encryption to members of a committee randomly composed of several clients. After the aggregator obtaining the total encrypted summation by public key from clients, it sent the result to the committee who then applied MPC to “fix” the secret the key and decrypted the result before revealing it. Functional encryption is another technique related to MPC. It expands upon public-key encryption by allowing an individual with a secret key to decipher a particular function derived from the encrypted content in the ciphertext. In [98], authors proposed the FAA-DL where clients and the server can collaboratively and proactively analyze anomaly based on functional encryption.
In Fig. 4, we demonstrate the mechanism of some privatization techniques. For the cryptography-based aggregation, LDP, and DDP, we take the sum computation task as an example.
V-B Analytics techniques
With the development of FA, there’s a paramount requirement for effective scalable, and private techniques to aggregate insights across decentralized clients. It’s communication-intensive to upload encrypted data sets and it’s insecure to expose data even with noise while some transformation of data structure may effectively address these challenges. In this section, we discuss the data sketching and some interesting specialized data structures that artfully assist FA.
Sketching. In FA, the pursuit of insight aggregation across decentralized data nodes demands techniques that are both efficient and scalable. Sketching stands out because it is a method designed with specific attributes tailored for the challenges posed by FA. These attributes [99] are query-specific (designed for specific insight extraction tasks), mergeable (allowing for easy aggregation from multiple clients), and extremely compact (ensuring minimal communication cost). Furthermore, various sketching methodologies have been developed to cater to specific data types and tasks. For instance, for set data, the Bloom filter is suitably employed for set intersection tasks. Similarly, the count sketch method is adept at identifying heavy hitters. To address the paramount privacy requirements in FA, sketching often collaborates with privatization methods such as noise addition, randomized responses, and hash functions to ensure the privacy and security of the aggregated insights. Together, these techniques present a robust and private solution for insight aggregation in FA.
Bloom filters [100] is a space-efficient probabilistic data structure used to test whether an element belongs to a set with possible false positives and no false negatives. Bloom filter and its variants are first introduced to federated analytics in Private Set Intersection tasks [101, 102, 103, 104]. Researchers in [101] combined Bloom filters with Goldwasser Micali homomorphic encryption to compute the intersection which is secure in malicious model. In [102], authors proposed a (PSI) method using Bloom filters with a secure multiplication protocol SEPIA to obtain an intersection of Bloom filters for each party to find their intersections. Authors in [103] reduced the hash operations to propose a faster protocol for the PSI task by integrating oblivious transfer with a garbled Bloom filter (GBF). The essential difference between the GBF and Bloom filters is that the GBF uses an array of $\lambda$-bit string where $\lambda$ is a security parameter while Bloom filters use an array of bits. Specifically, the client computes its sets to a Bloom filter and the server also computes its set to a GBF. Then an oblivious transfer protocol is conducted to give the client a GBF of intersection which is finally queried by the client to obtain the result. There were continuing works on Dong et al. [103] such as optimization of performance based on random OT extension[104], enhancement in its malicious-secure variant [105], and multi-parties intersection[96, 106]. RAPPOR [15] introduced by Google is a widely applicable and practical data collection mechanism providing strong privacy guarantees with high utility. The basic idea of RAPPOR is applying randomized responses to Bloom filters that each client reports each bit of her Bloom filter with possible untruthful responses to the server for privacy guarantees.
The counter, a common data sketching method, is a dictionary data structure storing the number of occurrences of elements. Apple developed two LDP algorithms based on it: Count Mean Sketch (CMS) and Hadamard Count Mean Sketch (HCMS) to learn dictionary-related tasks such as learning particular words and discovering popular emojis [82]. The main idea is to index counters sketched by raw data at the client side by hash functions and Hadamard transform with randomized responses. Researchers in [25] proposed TreeHist protocol to solve heavy hitters problem. The protocol utilized a local randomizer which could be regarded as a sampled and noisy version of the count sketch. The sketch is then Hadamard transformed to bits and one bit is sampled for subsequent submission to the server and aggregation to binary prefix tree. In [44], authors proposed methods that encode counters by Laplace noise addition without costly reconstruction like hash-based techniques under a central DP setting. They saved the computation resources of the aggregator by avoiding reconstructing results from perturbed messages.
TABLE V: Summary of sketching.
Reference Task Sketching method Cooperated privatization technique
Kerschbaum [101] PSI between 2 parties Bloom filters Goldwasser Micali Encryption
Many et al. [102] Multi-parties PSI Bloom filters SEPIA library
GBF [103] PSI between 2 parties Bloom filters Oblivious Transfer
Pinkas et al. [104] PSI between 2 parties Bloom filters Oblivious Transfer
Rindal et al. [105] PSI between 2 parties Bloom filters Oblivious Transfer&Choose-and-cut
Kolesnikov et al. [96] Multi-parties PSI Bloom filters Oblivious Transfer&Pseudorandom functions
Inbar et al. [106] Multi-parties PSI Bloom filters Oblivious Transfer
RAPPOR [15] Data federation Bloom filters Randomized responses
CMS&HCMS [82] Dictionary-related tasks Count sketching Hash functions&Hadamard transform
TreeHist [25] Heavy hitters Count sketching Hash functions&Hadamard transform
HH&PEM [44] Heavy hitters Count sketching Noise addition
Specialized data structures. In the context of FA where data analysis is performed across federated devices and servers, some specialized data structures are proposed to help the server to derive the analytics results on particular tasks. In [18], authors proposed the TrieHH algorithm discovering frequently typed words in edge devices, which is an iterative algorithm to find heavy hitters with the Prefix Tree. The Prefix tree, or the Trie, is a tree-like data structure used to store strings like words. Each node in the Trie represents a single character of a string and all end nodes represent an end identifier so that each end node stands for a string and all sibling nodes own a common prefix. In their TrieHH algorithm, a subset of clients is sampled in each iteration and the Trie is updated based on their data points after filtering by a threshold until convergence. This sampling-and-threshold algorithm is used to provide central differential privacy instead of the usual noise addition. There’s a following work of it, TrieHH++[43], which answered more general queries that not only identified heavy hitters but also provided their estimated frequency. In [107], researchers proposed OptPrefixTree based on previous work. They suggest a flexible algorithm to increase efficacy by utilizing adaptive segmentation, intelligent data selection, and deny lists.
There’s also a tree-based method utilized in [29] to guarantee robustness and prevent the aggregator from cheating, which is a summation tree. Authors proposed the Honeycrisp system which uses sparse vector theorem to schedule the DP budget. The aggregator only has ability to conduct summation on collected encrypted data. The committee composed of some clients can decrypt the result and compare it with the guess from the aggregator to decide whether to reveal it. When conducting the total summation in the aggregation phase, the aggregator is asked to generate a summation tree in which each parent node owns two children nodes and is the sum of them under homomorphic encryption. The tree is then checked by clients with each client verifying one pair of a parent and its children to prevent the aggregator from manipulating the aggregated result.
V-C Deployment and optimization techniques
Beyond the basic privacy afforded by restricting raw data within the data owner, there’s the complex challenge of deploying effective techniques for optimization in some scenarios. These techniques are essential to navigate the distinct problem FA presents in terms of computation, communication, data heterogeneity, incentive mechanism, utility, and privacy. In this section, we list some techniques concentrating on the deployment and optimization problems in FA.
Optimization theory. The optimization challenges in federated analytics differ significantly from those in a centralized setting. As mentioned in [108], federated optimization mainly focuses on communication efficiency, data heterogeneity, computation, and privacy constraints.
In [109], FedVision is introduced, optimizing resource usage in video analytics. This system minimizes network resource consumption and maximizes computational efficiency by integrating black-box optimization with Neural Processes, tailored for dynamic network conditions. The study in [110] presents Edge-DemLearn, an approach that maximizes learning performance and system efficiency. By leveraging distributed computing infrastructure, Edge-DemLearn is a sub-optimal two-sided many-to-one matching algorithm optimizing both the allocation of resources and the generalization of models. This method effectively overcomes the inherent limitations of traditional Federated Learning by enhancing model generalization and optimizing the management of user equipment associations within a distributed learning framework. Paper [111] represents a method that combines federated analytics with 5G technology to optimize data collection and processing in vehicular networks. This approach aims to minimize latency by employing a federated analytics-based solution and a D/M/1 queuing systems-based mathematical model to analyze end-to-end latency. They utilize numerical optimization methods like linear approximation (COBYLA), the Trust Region Constrained Algorithm (TRCA), and the Sequential Least SQuares Programming (SLSQP) to optimize system parameters, focusing on reducing waiting times and maximizing information value. S. Shi et al. in [112] introduced a Distributionally Robust Optimization (DRO) paradigm to optimize robustness in network traffic classifier learning under noisy labels. DRO is a modeling approach that makes decisions under uncertainty by minimizing the worst-case cost across all possible distributions in a constructed uncertainty set. This strategy maximizes classifier accuracy despite the variability in data quality.
Game theory and incentive design. In federated settings, incentive mechanisms are crucial for motivating participants to contribute their private data, communicational resources, and computational resources. Incentive design typically balances individual rewards with collective goals to guarantee that participants are fairly compensated and to maintain the efficiency of the system. By employing game theory approaches such as contract theory and multi-leader-follower games, these incentives aim to align participant actions with the overall objectives of the federated tasks. The structure of incentive mechanisms in FA is considerably similar to those in FL, where collaborative data processing and performing tasks require nearly the same strategies. Game theory in federated analytics emerges as a critical tool for resolving complex interactions and optimizing cooperative strategies across various domains. It efficiently addresses challenges in privacy and incentivizes participation, scalability, and effectiveness in decentralized systems.
In [113], researchers introduce CROWDFA which employs federated analytics in mobile crowdsensing and balances data aggregation, incentive design, and privacy preservation. It uses additive secret sharing for privacy and a novel incentive mechanism, PRAED, to encourage participants’ involvement with privacy protection. Authors in [114] integrate blockchain with federated learning in healthcare metaverses, focusing on a user-centric incentive mechanism. It introduces AoI as a metric for data freshness and employs contract theory to incentivize data sharing. The framework uses Prospect Theory to address the service provider’s decision-making under uncertainty, aiming to enhance immersive healthcare experiences by ensuring data freshness and privacy. Another incentive mechanism is developed in [115] for federated learning and analytics systems with multiple tasks. It uses a multi-leader-follower game where rewards are set to motivate data owners, who in turn decide their participation level. The paper provides algorithms for optimal strategy formulation, ensuring effective resource allocation and addressing the challenges of multi-task federated systems. S. Shi et al. in [116] focuses on updating HD maps for autonomous driving using a game-theoretical model. It proposes an overlapping coalition formation game where vehicles collaborate in coalitions to update map data for maximizing utility and map quality. It addresses privacy and incentivizes participation by allowing vehicles to join multiple coalitions, which benefits the scalability and accuracy of map updates. In [117], researchers cope with privacy and pricing in data markets. It proposes a federation model where data providers form coalitions with differential privacy. They introduce a method for determining collective data prices based on privacy levels and employ the Shapley value from game theory for fair earnings distribution, which benefits both data providers and consumers by balancing privacy concerns and financial incentives.
Figure 5: Utilizing sampling in federated analysis reduces computation and communication costs by reducing the amount of data processed. Although it introduces result variance due to analyzing less data, it can amplify privacy budgets to improve result accuracy by reducing data perturbation.
Sampling. To aggregate insights across these distributed clients, federated analytics often utilizes sampling techniques in terms of enhancing the privacy and improving the performance. There’s also an optimization-wise problem for sampling techniques and the accuracy as shown in figure 5. 1) When sampling is utilized, the total amount of data decreases so that the privacy budgets are amplified. The accuracy of results is improved since the federated data are less perturbed with more privacy budget. 2) While intuitively, less data analyzed leads to more variance of the results that sampling may cause accuracy reduction.
In early research within the domain of differential privacy, the sampling technique is referenced and employed for the purpose of privacy amplification[118]. Instead of examining the full data from all clients, using either data from a sample of clients or each client’s sampled subset of data can enhance privacy protection in terms of amplify the privacy bounds of a differentially private mechanism under various conditions. In other words, the utilization of sampling has ability to improve the accuracy with the ”amplified” privacy budgets. In most time, sampling is combined with other privatization techniques like the noise injection to further protect data. Researchers in [38] proposed the LDPMiner, a two phases mechanism for heavy hitters task which applied sampled randomizer (combining sampling with privatization techniques like RAPPOR and Succinct Histogram) to each phase. They conducted the experiments on synthetic data and one of results is the accuracy of targets from sampled randomizer outperforms that from naive randomizer under the same privacy budgets. It’s worth noting that there’s a born-in limitation of sampling that it’s always require large enough total data to ensure its effectiveness, which is also indicated by another experimental results in [38]. There’s a continuing work on LDPMiner, which proposed SVIM [41], a protocol for searching frequent items under set-valued LDP setting. They attempted new cooperated privatization techniques with sampling and found that some techniques benefit privacy amplification from sampling like GRR (Generalized Random Response) while some doesn’t such as OLH (Optimized Local Hashing). They explained the interesting phenomenon that the relation between reported and input values needs to satisfy a ”many-to-one-wise” mapping for benefit of privacy amplification from sampling, i.e., reported value after encryption ”support” one input value instead of multiple values [41]. Also, sampling can improve efficiency of system attribute to the reduction in amount of data transmitted and processed while facing with the concern of accuracy loss. In [49], authors proposed a private data federation system SAQE (Secure Approximate Query Evaluator) and denoted that sometimes sampling improves not only the efficiency but also the accuracy, i.e. utilizing less data in the computation leads to improvement in accuracy. Given invariant privacy budgets, the sampling error decreases as the sampling rate grows up, which means it’s required more noise injection to maintain the differential privacy. Next, the total error summed by sampling error and error from noise is a convex function of sampling rate for some tasks, meaning that sampling with proper rate contributes to an increasing accuracy. Now that sampling has the ability to introduce error to guarantee privacy, there are also attempts on merely applying sampling instead of as a bonus privacy amplifier with other techniques like noise injection.
TABLE VI: Summary of sampling.
Reference Sampled Object Sampling rate Sampling method Cooperated privatization technique
LDPMiner [38] Data of each client 1 over data number for each client Uniform Succint Histogram&RAPPOR
SVIM [41] Data of each client 1 over data number for each client Uniform GRR&OLH
SAQE [49] Data of each client Optimized result based on the task Bernoulli Noise addition
VI Statistical metrics
Statistical metrics, like the means, variances, medians, and percentiles, are simple but significant components in the field of data analytics. Deriving statistical metrics has various real-world applications, and also acts as intermediate steps of many complex data analytics algorithms. The demand for computing the statistical metrics of client data with privacy preservation motivates numerous FA studies. In this section, we focus on three widely-used FA applications on statistical metrics: mean (Section VI-A), median/percentile (Section VI-B), and clustering center (Section VI-C).
VI-A Means
The FA studies on calculating means usually consider that each client possesses a scalar value [^3] The case of vector values can be formulated by executing multiple FA tasks on scaler values independently., i.e., consider totally $n$ clients, where client $i$ possesses a scaler value $d_{i}$. The FA system derives the mean of their client-held scalers:
$\mu(d):=\frac{1}{n}\sum_{i=1}^{n}d_{i}.$ (6)
Mean computation is likely to be the most widely used data analytics task. It can be also transformed into other metrics. For example, sum computation can be conducted by multiplying the mean by $n$, and variance computation $V(d)$ can be derived as follows by executing the mean computation for the client data $d_{i}$ and square of client data $d_{i}^{2}$.
$V(d)=\mu(d^{2})-\mu^{2}(d)=\frac{1}{n}\sum_{i=1}^{n}d_{i}^{2}-\big{(}\frac{1}{n}\sum_{i=1}^{n}d_{i}\big{)}^{2}.$ (7)
As a strong and formal guarantee, LDP is applied in many FA-based mean computation solutions. LDP can be straightforwardly applied on mean computation by perturbing the client data $d_{i}$ before being uploaded. Based on the simple idea, researchers designed various solutions with their unique advantages. Cormode and Markov [21] propose an FA solution for the mean computation. In addition to satisfying the strong privacy criterion of LDP, their solution has a unique advantage, in that each client in their solution only needs to upload one bit to the server. Although it does not improve privacy preservation from a mathematical perspective, uploading only one bit is convincing for users with little mathematical backgrounds that the exposed information is minimized. In [22], an LDP solution is proposed to privately estimate the mean and histogram of distributed data. It also proposes a discretization technique to handle the continuous collection of user data (e.g., user data are collected daily). Harmony [23] provides an LDP solution for data analytics in mobile phone applications. It realizes mean estimation based on the Laplace mechanism, and frequency estimation based on the randomized response. PrivKV and its variations [24] extend the statistic tasks of mean estimation into the setting of data in the key-value structure. In that setting, each client holds a key-value pair; the data analyst expects to sort the keys of the clients and calculate the means of values with each same key. This paper then proposes an iterative structure to realize LDP with improved data utility. Honeycrisp [29] is a system designed to address large-scale count mean sketch tasks without a trusted server. A committee is randomly selected from users to generate homomorphic encryption keys. Client data is sent to the server using additively homomorphic encryption, which allows the server to calculate the total sum without accessing individual raw data. This sum is then decrypted by the committee. It utilizes the sparse vector theorem to manage the DP budget, enabling the committee to decrypt results and verify the aggregator’s calculations before deciding on their disclosure. In this case, the system also includes a summation tree method to ensure robustness and prevent the aggregator from cheating, where each node is the sum of its children under homomorphic encryption and they are verified by clients to ensure integrity.
In the aggregation phase of FL, the FedAvg algorithm is computed on the client uploads. Essentially, the mean of gradient vectors from different clients is computed as the update on the global model. Many FL researchers try to improve the privacy and security of the aggregation phase, which falls into the category of FA-based mean computation.
Figure 6: A minimal running example of pairwise mask-based secure aggregation.
Secure aggregation-based approaches aim at improving the privacy preservation of uploading gradient vectors via cryptography tools. With secure aggregation, the server can only learn the aggregated result (mean of gradient vectors), but cannot obtain any knowledge about the uploads of any single client. Pairwise masking is the most lightweight and easy-to-implement technique to realize secure aggregation. In pairwise masking, each client communicates with some other clients, and a pairwise mask is generated for each pair of clients. The pairwise mask is added to the upload in one client in the pair, and its inverse is added to the other client in that pair. When an aggregator sums up the client uploads, the pairs of masks are eliminated, and the sum of the original data can be revealed. Fig. 6 provides an example of pairwise masking. In Fig. 6, three clients, holding local upload $X_{1},X_{2}$, and $X_{3}$ respectively, form three pairs and generate three pairwise masks. For example, mask $m_{1,2}$ is generated for clients 1 and 2, which is added to client 1’s upload, and inversely added to client 2’s upload. By summing up masked client uploads $U_{1},U_{2}$, and $U_{3}$, the sum of $X_{1},X_{2}$, and $X_{3}$ is revealed.
In [25], a pioneering secure aggregation scheme for FL SecAgg is proposed. It leverages the agreed random seed of client pairs to generate random masks on the uploads. The clients also apply self-masks on their uploads and leverage secret sharing to propagate the information of their self-masks. With the utilization of two kinds of masks, the FL system can preserve functionality and security with the existence of one-third of malicious clients and another one-third of dropout clients. In [26], the authors consider the heavy computation and communication loads of deriving and eliminating the masks and propose an improvement on SecAgg. In the new design, each client no longer needs to communicate with all other clients to share the masks. Instead, a client only needs to interact with a small random part of clients, which are logarithmic to the client size. The new design is still able to defend one-third of malicious clients and another one-third of dropout clients with high confidence, but not with 100% confidence. In [27], FastSecAgg is proposed to improve the vanilla SecAgg protocol. It relies on a novel multi-secret sharing scheme based on fast Fourier transform, so that the per-client workload is reduced, but the guarantee on the portion of defended malicious/dropout clients is weakened. In [28], Turbo-Aggregate is proposed that employs the circular communication topology to reduce the communication overhead, but sacrifices privacy preservation as it only guarantees privacy in the average case, instead of the worst case considered in SecAgg.
Some other studies replace pairwise mask generation, which requires extensive client pair interactions, with non-pairwise masks. In non-pairwise mask generation, each client generates masks independently, and the server can eliminate the masks by learning the sum of the masks with a special protocol. In [119], a trusted third party is employed to assist with the masking, unmasking, and dropout handling. In LightSecAgg [120], the previous protocol is improved by removing the need for a trusted third party, based on a secure and dropout-resilience secret sharing scheme. In [121], the homomorphic pseudorandom generator (HPRG) is leveraged as a non-pairwise mask generation scheme to achieve lightweight, dropout-resilience secure aggregation.
DP-based approaches are also employed in privacy-preserving gradient aggregation. LDP is widely used in these solutions for its strong privacy guarantee. In [83], user-level DP is satisfied by applying Gaussian noise on the gradient vectors to enforce LDP. The proposed approach also provides a theoretical FL convergence guarantee, and trade-off between convergence speed and computation overhead. In [84], another LDP-enhanced aggregation approach is proposed by applying a Gaussian mechanism to the gradients. [84] also provides privacy-utility-communication trade-off, adjustable query sensitivity for LDP, and an accounting scheme of LDP budgets. LDP-FedSGD algorithm [84] considers privacy-preserving FL aggregation for IoT applications. It realizes LDP by perturbing the gradients with four LDP mechanisms. These mechanisms have different output structures and are optimal for different ranges of $\epsilon$ values. In HFL-DP [122], FL under the server-edge-client architecture is considered. The employed Gaussian mechanism provides an LDP guarantee under different levels on client uploads and edge uploads. Since the FL, and also the aggregation phases, take many rounds to complete, and the LDP budgets are cumulated by the composition theorem, naïvely applying LDP in each round could result in quite a low privacy budget for each round. Therefore, many of the aforementioned works [83, 84, 122] propose additional privacy-utility trade-offs or privacy budgeting schemes to tackle the issue.
On the other hand, since LDP typically introduces significant noise on the uploads, which harms the accuracy of mean computation results, some researchers investigate privacy-preserving gradient aggregation approaches satisfying the CDP criterion. In [87], the authors consider applying CDP on the aggregated gradient, based on a trusted aggregator (server). Such a CDP-enforced gradient can hide the participation of each participating client so that the privacy of each client is preserved. Gaussian noise is adopted to enforce CDP in [87]. In [88], another privacy-preserving aggregation scheme where similar settings of trusted aggregator, client-level privacy, and Gaussian mechanism, are considered. It is particularly designed for recurrent language models. The idea of CDP-enhanced aggregation is also employed in variations of FL like personalized FL [86], applying the Gaussian mechanism.
Privacy-preserving aggregation for FL, or privacy-preserving vector mean computation from the FA perspective, has gained numerous research interests in recent years. A thorough survey of these researches is in [123].
VI-B Medians and percentiles
Median and percentile computation in the federated setting considers the clients each with a number of scaler data points $d_{i}$. $\mathcal{D}=\{d_{i}|i\}$ is the virtually aggregated data points from all clients. Consider a ranking function $\mathcal{R}(\mathcal{D},k)$ outputting the element in $\mathcal{D}$ exactly larger than $k$ portion of the elements in $\mathcal{D}$. The federated $k$-percentile computation lets the clients collaboratively derive the value of $\mathcal{R}(\mathcal{D},k)$, while preserving the privacy of data points $d_{i}$. The federated median computation is a special case of the federated $k$-percentile computation. The major challenge of federated median/percentile computation is that it requires different clients to compare their private data points to derive the ranking results.
As the private scaler computation is a quite classical problem setting in the field of MPC, some researchers design federated median/percentile solutions utilizing various cryptography tools. In [30], a private solution of median (and $k$-precentile) element over client local datasets is proposed. The solution is based on SMC and is realized by transforming the problem into a combinatorial circuit. In [31], a federated median/percentile computation scheme is proposed following the server-client architecture. The clients provide privatized input to the computation, and the server conducts the computation-intensive workloads to derive the results without observing the raw data. LDP is also employed as a formal privacy guarantee in federated median/percentile solutions. In [89], a median calculation scheme is proposed with both LDP and MPC. It can be extended to tasks like percentile computation. The LDP guarantee is realized by an exponential mechanism. The authors consider threat models of both semi-honest settings and adversarial settings. Another federated median/percentile computation solution [32] relies on a model-based optimization approach in a federated way, without a formal privacy guarantee. In [32] deriving the quantile is formulated as an optimization problem. That optimization problem is then collaboratively computed by the clients with ADMM. No formal privacy guarantee is provided.
VI-C Clustering
The clustering task aims at dividing data points (usually defined by fixed-length vectors) into groups, so that data points within a group are more similar to each other in some sense (e.g., with lower vector distance), and data points in different groups are less similar. Clustering analysis is a major task in the field of exploratory data analytics and is an important technique in the fields of pattern recognition, sociology, data compression, and bioinformatics.
The federated clustering task extends the clustering problem into the federated data setting. In that setting, the clients, each possessing a subset of data points, want to collaboratively derive the global clustering results, i.e., deriving what groups their data points belong to after the data points from different clients are virtually aggregated. The major challenge of federated clustering is that it needs to perform similarity measurements for data points from different clients while preserving data privacy.
Researchers have proposed federated clustering solutions following the FA paradigm. In these proposed algorithms, classical centralized clustering algorithms are adapted into the federated version, with the intermediate results serving as the FA insight. [17] proposes a federated clustering solution. It utilizes a hierarchical structure, where data points are first clustered locally within local client data, and local centroids are then clustered by the server. The authors derive an interesting finding, that in contrast to other federated computation solutions, data heterogeneity is beneficial to the analytics results. In [33], the authors consider kernel $k$-means, an effective clustering algorithm in capturing the nonlinear representation of the dataset using the nonlinear kernel functions, in the federated setting. To realize that, the clients first transform their data points into feature vectors by applying the kernel function. UIFCA [34] considers the unsupervised clustering tasks (the authors take unlabelled image clustering tasks as an example). It follows the idea of the IFCA algorithm, which trains a model for each cluster, and assigns a data sample to a cluster with the smallest model loss. To modify IFCA, which was originally designed for personalized FA, UIFCA replaces the models used to be generative models that capture the distribution of one cluster.
In [35], a self-supervised learning-based federated clustering algorithm Orchestra is proposed. It takes a Sinkhorn-Knopp-based deep learning clustering algorithm [124] which outputs equal-size clusters. The clients first cluster their local data to derive local centroids, and upload both the cluster model parameters and the local centroids to the server. In the server, the model parameters are aggregated by FL, and the local centroids are aggregated by FA via applying the Sinkhorn-Knopp algorithm again. Then, the updated model parameters and global centroids are sent back to the clients to start the next round, until a desirable clustering result is achieved. In [36], the FA technique is employed as a prepossessing step of a federated clustering algorithm. In that FA design, the data space is divided into a grid, and each client encodes their local data into a vector, representing the number of samples falling into each bin of the grid. Then, the vector is uploaded to the server as insight. The server aggregates the insights from all clients and derives a global view of data distribution. After that, the grid is updated following the method of the self-organizing map, and the updated map is sent to the clients, where the clients start a new round of bin checking.
VII Frequency-related applications
The tasks related to the frequency estimation on categorial data make up an important part of data analytics, and also its FA variations. In these tasks, the clients with indexes $i=1,...,n$ hold local data $d_{i}$. Denote $q$ as a queried data structure, for any $q$ and $d_{i}$, there exists a relationship of “containing”. For simplicity, we borrow the symbols from set theory, so that $q\in d_{i}$ denotes the fact that “$d_{i}$ contains $q$”, and $q\notin d_{i}$ denotes the fact that “$d_{i}$ does not contain $q$”. For any $q$ and $d_{i}$, exactly one of $q\in d_{i}$ or $q\notin d_{i}$ must be true.
With the definition of a “containing” relationship, the frequency $f$ of any query data can be then defined as follows.
$f(q)=\frac{|\{i|q\in d_{i}\}|}{n},$ (8)
or
$f(q)=\mathbb{P}(q\in d_{i}),{\rm~{}randomly~{}sample~{}}i{\rm~{}from~{}clients}$ (9)
can be used when the size of clients $n$ is not available.
With the definition of the frequency of queried data, multiple frequency-related tasks can be then defined. Here we list our taxonomy of these tasks.
Frequency oracle. The task of frequency oracle aims at constructing an abstraction (frequency oracle) of client data. Using the abstraction, a querier can arbitrarily derive estimations of any queried data without interacting with the clients. In some works, the frequency oracle problem is termed frequency estimation problem as the frequency oracle provides services of frequency estimation.
Frequent pattern mining. The task of frequent pattern mining (FPM) aims at finding highly frequent queried data. There are two mainstream definitions of FPM. In the first definition, the system outputs all possible queried data $q$ where $f(q)$ is higher than a predefined threshold. In the second definition, the system outputs the top-$k$ queried data with the highest frequency. Based on the structures of client data $d_{i}$, queried data $q$, and definition of containing relationship, FPM is also named by its subtasks frequent item mining, frequent itemset mining, and frequent sequence mining.
Heavy hitter. The term heavy hitters originated from the downstream tasks in web data analysis. The heavy hitter task is consistent with frequent item mining, where the containing relationship is defined to be the $\in$” relationship in set theory.
The definitions of frequency-related tasks are highly connected, and many algorithms are able to tackle multiple tasks simultaneously.
In the following part, we will first discuss some key issues in handling frequency-related tasks in Section VII-A, and then introduce the existing FA solutions in Section VII-B
VII-A Key issues in frequency-related tasks
Data domain and data type are two key issues in the problem setting of any frequency-related task, which determine the difficulty of the task and introduce challenges in designing the solutions.
Data domain. The data domain (or data universe) is an important setting in frequency-related tasks. The data domain refers to the set of all possible queried data. The easiest setting is the small domain, where the number of possible queried data is small enough so that the algorithm designer can simply enumerate all possible queried data, or design one-hot encoding of them. In the large domain setting, the algorithm designer still has knowledge of all the possible queries, but can no longer apply one-hot encoding or similar techniques, due to the limit of computation and communication resources. In the most difficult infinite domain setting, there are infinite possible queried data. An example of the infinite setting is that the queried data are all possible strings of characters with arbitrary lengths.
Data type. The frequency-related tasks can be categorized by the type of queried data (and also client data). The item type data is the default setting of frequency oracle and heavy hitters, where the possible queried data are indivisible distinct items. Complex types of queried data, like set, sequence, and graph of items, are investigated in frequent pattern mining problems. These problems are typically more difficult than the item-type problems, because they lead to exponentially increasing data domain, and require careful design to handle the relationship between different queried data.
VII-B Existing solutions
The majority of FA-based frequency oracle, heavy hitters, and frequent pattern mining solutions adopt LDP for their privacy guarantee, owing to its strong privacy preservation in the untrusted aggregator model. RAPPOR [15] is one of the earliest and most popular frequency oracle solutions with LDP guarantee. It lets the clients encode their local data into bloom filters, so that LDP can be enforced for arbitrary structure of local data. By aggregating the bloom filters, the frequency of the encoded elements can be estimated. [37] (succinct histograms) is another algorithm for frequency oracle and heavy hitter problems. With the help of shared randomness, the succinct histogram algorithm only requests clients to upload one-bit data. LDPMiner [38] is another solution for heavy hitters. Compared to previous solutions, LDPMiner is the first one to resolve the set-value client data, where multiple $q$ can be contained by one $d_{i}$. PSFO [41] proposes its optimization over LDPMiner. While preserving the strengths of LDPMiner, PSFO enables frequent itemset mining tasks with its padding-and-sampling oracle. In [42], two algorithms, TreeHist and Bitstogram, are proposed, to resolve the tasks of frequency oracle, and heavy hitters. In the two algorithms, tailored data structures, shared randomness, and Hadamard transform are carefully designed and utilized. As a result, TreeHist and Bitstrogram manage to optimize the time and space complexity on both the server and client side. In [39], another frequency oracle algorithm based on Hadamard response is proposed that removes the need for shared randomness. In [40], an algorithm is proposed for frequency oracle and heavy hitters. In addition to removing the need for public randomness, [40] only requires one-bit responses. FIML [125] tackles the top-$k$ setting of frequent itemset mining. It first builds a frequency oracle to find out top frequent items, then builds candidate itemsets, and queries their frequencies. OptPrefixTree [107] is a prefix tree-based heavy hitter algorithm. It leverages the prefix tree to handle the infinite data domain.
SFP [82] algorithm is proposed by Apple to tackle the difficult frequent sequence mining problem. It encodes the local client data with count mean sketch so that LDP can be properly enforced. It partially tackles the large domain problem of domain element frequency estimation. It lets clients upload an extra count mean sketch of sequence fragments, where the short fragments can be easily enumerated and frequent short fragments can assemble longer sequences. FedFPM [16] is proposed as a unified FA framework to tackle multiple FPM subproblems, including heavy hitters, frequent itemset mining, and frequent sequence mining. It follows a query-response scheme to estimate the frequency of the candidate patterns and derives Hoeffding’s inequality-based bounds to filter the candidates. FedFPM is able to achieve a better data utility than existing solutions but requires sufficient participating clients.
As LDP schemes request clients to add significant noise to the uploads, researchers propose CDP schemes in FA-based frequency-related tasks. The major advantage of CDP is that it requests clients to add smaller noise on the uploads, or even add no noise by utilizing the sampling mechanism. However, only enforcing CDP sacrifices the formal privacy guarantee, as the FA server is usually considered untrusted. TrieHH [18] is an FA solution to digest string-typed heavy hitters. It iteratively builds up a prefix tree of characters with the interaction between the server and clients. Its major advantage is the ability to satisfy CDP without adding random noise to the outputs so that a better data utility of the mined results can be achieved. TrieHH++ [43] is an expansion of TrieHH. It extends the applications from heavy hitters to quantile estimation and range query. It follows the basic ideas of TrieHH, that realizing CDP via sampling data without adding random noise to the outputs and generating a prefix tree to handle infinite data domain. In addition, it applies Poisson sampling, instead of fixed-size sampling, to hide the sample size.
Some works adopt cryptography tools instead of DP in FA-based frequency-related solutions. Compared to DP, cryptography tools allow deriving accurate results with maximal data utility. However, an adversary still may infer sensitive information from the accurate outputs. STAR [20] is a threshold aggregation (a variation of heavy hitters) solution for web data analysis. Its threshold aggregation enforces $k$-anonymity on the output data, and such functionality is realized by a secret sharing scheme. In [126], a private heavy hitter algorithm is proposed. The solution relies on the incremental distributed point function, a lightweight cryptography tool. It assumes that there exist two servers that collude with neither each other nor any client. In [44], a private heavy hitter solution is proposed that utilizes DDP for its privacy preservation. DDP statistically satisfies CDP, and utilizes cryptography tools to prevent the server from learning any single client upload, so that local privacy can be achieved with CDP-level data utility. In FedWeb [45], another DDP-based solution is proposed. It can handle generic FPM tasks, and proposes tailored design to be applied in Web 3.0 applications.
In [12], an FA-based location heatmap generation solution is proposed, which can be regarded as a frequency oracle solution on geometric items. In [12], DDP, which combines the schemes of cryptography tools and DP, is applied to reduce the noise required for privacy preservation. It iteratively reduces the granularity of queried locations to better utilize the participating clients.
In Table VII, we summarize our surveyed FA-based solutions on frequency-related tasks and list their detailed tasks, privacy model, data domain setting, and extra features.
TABLE VII: Summery of FA-based solutions on frequency-related applications.
Reference Task Privacy Data domain Notes
RAPPOR [15] FO & HH LDP Large Randomized response on modified bloom filter
Bassily et al. [37] FO & HH LDP Small One-bit client upload
LDPMiner [38] FO & HH & FSM LDP Small Constant communication cost
PSFO [41] FIsM LDP Small Converting FIsM into HH via padding and sampling on itemsets
TreeHist [42] FO LDP Small Near-optimal error
Bitstogram [42] FO & HH LDP Small Near-optimal error
Acharya et al. [39] FO LDP Small Hadamard Response; No shared randomness
Acharya et al. [40] FO& HH LDP Small Hadamard Response; No shared randomness; One-bit client upload
FIML [125] FIsM LDP Small Top-$k$ setting of FIsM
OptPrefixTree [107] HH & FSM LDP Infinite Handling infinite data domain with prefix tree
SFP [82] FSM LDP Large Uploading two count-mean-sketches to encode a sequence
FedFPM [16] HH & FIsM & FSM LDP Small Unified framework for multiple FPM subtasks; One-bit client upload
TrieHH HH & FSM CDP Infinite Building up the frequent strings by transmitting a prefix tree
TrieHH++ [43] HH & FSM CDP Infinite Randomized sample size
STAR [20] HH Cryptography Infinite Private threshold aggregation via secret sharing
Boneh et al. [126] FO & HH & FSM Cryptography Infinite High security via incremental distributed point functions
Böhler et al. [44] HH DDP Small Utilizing DDP for high data utility and formal DP guarantee
Bagdasaryan et al. [12] FO DDP Small Utilizing DDP for high data utility and formal DP guarantee
FedWeb [45] HH & FIsM & FSM DDP Small Application in Web 3.0 scenario
Figure 7: Illustration of the server-client-based federated query architecture.
VIII Database operations
Federated query, also termed data federation SQL, targets enabling database operations (SQL queries) on the data separated in multiple clients. The result of federated query is expected to be the same as the case where data from clients is virtually gathered to form a centralized database, and the privacy of the client data is preserved. Privacy preservation of federated query solutions is usually achieved by various cryptography tools so that every participant cannot learn any information about the data of other clients unless it is necessary to be revealed by the query result. Federated query has the potential to be an important piece of FA research, as many simple data analytics tasks are equivalent to a single SQL query, and some complex data analytics tasks can be decomposed into a series of database operations.
As privacy-preserving federated query solutions rely on cryptography tools, especially secure multi-party computation (MPC) protocols, the heavy computation overhead of executing the cryptography tools becomes a significant burden to deploy these federated query solutions into real-world systems. Therefore, an important task of the federated query solutions is to optimize the computation time of executing the federated query and deriving the accurate result. To fulfill that goal, researchers propose various solutions to pursue desirable computation time performance, which utilizes various cryptography tools with different properties and carefully optimizes the computation graph of executing the query to reduce the computation cost.
Federated query solutions usually follow two kinds of architectures. The first architecture is inherited from classical MPC protocols like garbled circuits. In such a fully decentralized architecture, clients directly communicate with each other under the encrypted protocol. After the computation, all the clients can derive the decrypted query results. Another widely applied architecture is the server-client architecture, as is illustrated in Fig. 7. The server is in charge of receiving the SQL statement from the federated query user and composing the computation plan. The clients execute the computation plan to conduct local computation, where client-client communication is usually involved as is required by the encryption tools. The clients upload the encrypted results to the server. Then, the server derives the plaintext SQL result by decrypting the client uploads, and returns the result to the federated query user.
Among the existing federated query solutions, the most ambitious ones are those supporting many types of queries (even the whole SQL standard) within one framework. These solutions are the most powerful FA solutions from the perspective of the number of supported data analytics tasks. SMCQL [46] is a pioneering federated query solution. It utilizes the garbled circuit to encrypt the whole query execution. The utilization of the garbled circuit makes SMCQL have the strong capacity to support the whole SQL standard but introduces significant computation overhead and only supports the computation over two clients. SMCQL also introduces a query planner to optimize the computation graph, so that some computation steps can be performed within one client and do not need SMC execution. Conclave [47] is then proposed to reduce the heavy computation workload of SMCQL. It utilizes the respectively lightweight secret sharing schemes to replace garbled circuit so that the computation is faster; computation among three clients is supported; but some query (e.g., window aggregate) is no longer supported. In addition, Conclave enables clients to optionally annotate some parts (columns) of their data to be non-private. The optional annotation reduces the scale of computation needed for SMC execution and further optimizes the computation time. Senate [50] is another SQL executor over federated data. Compared to SMCQL and Conclave, it changes the semi-honest assumption into a stricter adversarial threat model. It proposes novel SMC decomposition and query planning techniques, letting some computations be executed on a subset of clients. Saqe [49] considers a new direction of optimizing federated query runtime. It removes the restriction that the SQL result must be accurate, and let the federated query derive approximate results. As the execution over a small sampled set of data is enough to derive the query result with a bounded accuracy guarantee, Saqe achieves much better performance by reducing the size of the data to be computed. In addition, the data sampling scheme of Sage realizes an additional DP guarantee of the approximate query result. Orchard [53] is a novel federated query solution that transforms queries into three “zones” that computations on private individual data, aggregated but not noised data, and noised data, respectively. The results of local computation in first “zone” are homomorphically encrypted and sent to the aggregator which performs summation. The aggregated results are decrypted, noised, and announced by a committee later which owns the key. The detailed procedure is the same as its previous work [29] which is introduced in VI-A herein. Orchard can efficiently answer queries at scale as long as there is additive aggregation in them such as k-means, logistic regression, perceptron, and PCA. Arboretum [55] is designed to efficiently answer a wide range of queries in large-scale FA setups with potentially billions of participants. Arboretum’s key strengths include its ability to automatically optimize query plans and distribute computational tasks across participant devices. The system outperforms previous solutions by supporting new types of queries and matches the cost of existing systems that were hand-optimized for specific queries. Arboretum’s approach significantly enhances the scalability and feasibility of executing complex queries in distributed environments while maintaining strong privacy guarantees.
Although the aforementioned general federated query schemes obtain remarkable achievement in computation time reduction and client size scalability, the intrinsic difficulty of the general federated query still burdens its application in many runtime-sensitive and large-scale systems. As an attempt to compromise the types of queries, specific federated queries are proposed. Specific federated query solutions tackle one kind of query so that many tools other than general cryptography solutions can be applied. In other words, researchers only need to provide a cryptography version of a specific algorithm to enable a specific federated query. As a result, the specific federated query has the potential to achieve better performance in computation load and has wide application in scenarios where only a specific kind of query is required. In [48], a federated query solution is proposed to handle the free-connex join-aggregate query (a special kind of query). The authors provide a secure version of the Yannakakis algorithm over two clients, which is much more efficient in runtime compared to the garbled circuit. In [51], the authors consider skyline queries over vertical data federations. Their algorithm design decomposes the execution into local computation and cross-client secure aggregation, where the secure part is reformulated into the private set computation problem, which can be efficiently handled by MPC solutions. Hu-fu [52] tackles SQL queries related to spatial data. The authors notice that various spatial queries, (including kNN, range counting, and range query) can be realized by only three secure operations (summation, comparison, and set union), while other distance-based operations can all be conducted without encryption. Hu-fu then designs effective dedicated algorithms to handle each of the three secure operators and therefore achieves good performance in computation load and gains capacity in supporting up to ten clients. Mycelium [54] is a system that enables processing differentially private queries over large-scale distributed graphs, which are common in scenarios like disease or malware tracking. Mycelium achieves privacy by combining homomorphic encryption , a verifiable secret redistribution scheme (similar to Orchard [53]), and a mix network based on telescoping circuits. It can handle various queries relevant to medical research without compromising individual privacy or learning the graph’s topology.
In Table VIII, we summarize our surveyed federated query solutions, regarding their supported query types, privacy models, cryptography techniques, and maximal supported clients.
TABLE VIII: Summery of federated query solutions.
Reference Query type Privacy Cryptography technique Maximal client
SMCQL [46] General Encryption Garbled circuit 2
Conclave [47] General Encryption Secret sharing 3
Senate [50] General Encryption (adversary model) Garbled circuit 16
Saqe [49] General Encryption & DP Secret sharing 2
Orchard [53] General Encryption & DP Homomorphic encryption $>$1000
Arboretum [55] General Encryption & DP Homomorphic encryption $>$1000
Secure Yannakakis [48] Free-connex join-aggregate query Encryption Private set intersection (based on garbled circuit), secret sharing, & oblivious extended permutation 2
Zhang et al. [51] Skyline query Encryption Private set intersection (based on homomorphic encryption and garbled bloom filter) 10
Hu-fu [52] Spatial query Encryption Secret sharing 10
Mycelium [54] Graph query Encryption & DP Homomorphic encryption & onion routing $>$1000
IX Model-based applications
The model-based formulations and algorithms have gained great success in handling complex data analytics tasks. FA-based methods become an effective solution for model-based applications when the privacy-sensitive data stored in distributed clients are required for training the model. In this section, we review several well-developed model-based applications and their existing FA solutions, including linear regression model (Section IX-A), Bayesian model (Section IX-B), matrix factorization model (Section IX-C), multi-armed bandits model (Section IX-D), and unsupervised representation learning model (Section IX-E).
IX-A Linear regression
Linear regression is a classical model-based data analytics problem. It discovers the linear relationship between features and observations. Consider a dataset $\mathcal{N}$ with $n$ data points, for a data point with index $i$, $X_{i}$ is its vector in the form of a $k$-length vector, and $y_{i}$ is its observation in the form of a scaler. Denote $\boldsymbol{X}=[X_{1},...,X_{n}]$, and $\boldsymbol{y}=[y_{1},...,y_{n}]$, the model of linear regression assumes the linear relationship between $X_{i}$ and $y$, i.e.,
$\boldsymbol{y}=\boldsymbol{\beta X}+\boldsymbol{\epsilon},$ (10)
where $\boldsymbol{\beta}$ is the parameters of the linear regression model in the form of a $k$-length vector, and $\boldsymbol{\epsilon}$ denotes the independent noise for each data point. The objective of the linear regression problem is to derive a set of parameters $\boldsymbol{\beta}$ that can fit the relationship well, e.g., that minimize the square error (MSE), i.e.,
$\min_{\boldsymbol{\beta}}\sum_{i=1}^{n}(\boldsymbol{\beta X}_{i}-y_{i})^{2},$ (11)
or minimize the regularized square error, well known as ridge regression, i.e.,
$\min_{\boldsymbol{\beta}}\sum_{i=1}^{n}(\boldsymbol{\beta X}_{i}-y_{i})^{2}+||\boldsymbol{\beta}||^{2}$ (12)
In the federated setting, the whole dataset $\mathcal{N}=\{\boldsymbol{X},\boldsymbol{y}\}$ is distributed among multiple clients. For a better explanation, we construct $\mathcal{N}$ to be a matrix with $n$ rows and $k+1$ columns, and the $i$-th column is constructed by the $k$ entries of $X_{i}$, followed by $y_{i}$, so that each row of the matrix $\mathcal{N}$ exactly represents all information of one data point. Federated linear regression can be categorized into horizontal federated linear regression and vertical federated linear regression, based on how the data $\mathcal{N}$ is distributed among clients. In the horizontal setting, each client possesses a subset of rows of $\mathcal{N}$, i.e., each client possesses the complete information of a part of the data points. In the vertical setting, each client possesses a subset of columns of $\mathcal{N}$, i.e., each client possesses incomplete information (some dimensions) of all the data points. The two settings of federated linear regression introduce different challenges, resulting in varying techniques utilized.
In [127], the authors propose a horizontal federated linear regression solution, termed federated data analytics (FDA). It utilizes the fact that in horizontal federated linear regression, as a client possesses the complete information of some data points, it can independently evaluate the error of these data points, and derive the gradient of the linear regression parameters $\boldsymbol{\beta}$. FDA is based on a hierarchical architecture, i.e., the client’s local sets of linear regression parameters (low hierarchy) and a set of shared parameters held by the server (high hierarchy) are all updated based on the information learned from each other. The authors then provide adaptations of the linear regression model on various downstream tasks, including hypothesis testing, uncertainty quantification, and variable selection. In [128], another horizontal federated linear regression solution LDP-IHT is proposed, also applying the idea of gradient transmission. It applies the LDP scheme on the uploaded gradient for better privacy preservation. The authors then investigate the harm to data utility (represented by error bound) of applying LDP on the high-dimensional gradients and propose the sequential interactive algorithm LDP-IHT to reduce the harm. In [129], another solution is proposed, where the computation of horizontal federated linear regression is encrypted by homomorphic encryption and garbled circuit. This work takes a stricter setting, that one client only possesses features and observation of a single data point. DP-CFL [130] is a horizontal federated linear regression solution that transmits the coded version of the local features and observations to the server for privacy preservation. The coding procedure is designed to multiply the data matrix with a random matrix sampled from the Gaussian distribution so that DP guarantee is enforced by DP-CFL. The aggregated coded data is the intermediate result to compute gradient on $\boldsymbol{\beta}$, so the server can learn the optimized $\boldsymbol{\beta}$ based on the aggregated coded data.
In the vertical federated linear regression setting, an extra challenge is introduced, as a client can no longer evaluate any single data point, and clients need to collaborate to derive the full information of any data point with privacy preservation. In [131], an early vertical federated linear regression solution is proposed, where Powell’s algorithm is applied to let clients optimize their parts of $\boldsymbol{\beta}$, and a secure summation protocol encrypts the intermediate computation results. In [132], the authors propose an improved solution for vertical federated linear regression. They design a novel MPC protocol combining the garbled circuit and a tailored protocol for secure computation of the inner product. They propose a conjugate gradient descent-based solver for linear regression problems, which is more suitable for MPC schemes. Another vertical federated linear regression solution, proposed in [133], gains a major advantage: it only utilizes linear homomorphic encryption to achieve security, with the absence of other computation-intensive cryptography tools. In [134], the authors investigate the heavy communication overhead of the feature and observation sharing among clients and introduce the idea of performing computation on coreset, a small subset of data points that can derive an accurate enough linear regression result, instead of the whole dataset. The authors then propose an important sampling method to derive the coreset. The coreset solution is also applied in the federated clustering problem, which is introduced in the corresponding part of this survey.
In [135], the power of edge computing is leveraged to accelerate the vertical federated linear regression computation, where an erasure code-based scheme is utilized to allocate the client workload to the edge workers, and homomorphic encryption is utilized for privacy guarantee.
IX-B Bayesian model
Bayesian analysis is a powerful probabilistic tool to handle the prior and posterior distributions of complex probabilistic events. It has many important downstream topics. In Bayesian optimization, the characteristics of an unknown function are evaluated based on a set of observations (data points), while the function itself is considered a black box during the procedure. In Bayesian network learning, the casual relationship among a number of random variables is evaluated based on the observations. The FA version of Bayesian analysis is naturally motivated as the scenario is considered that the observations, necessary for any Bayesian analysis task, are held by different clients. The FA-based Bayesian analysis problem targets collaboratively deriving the result of Bayesian analysis with the observations held by the clients while preserving the privacy of the observations.
In [136], the authors propose a federated Bayesian optimization solution. It designs an FA version of the Thompson sampling to resolve the federated Bayesian optimization problem and proposes the random Fourier feature that enables the sharing of useful information for Bayesian optimization without transmitting raw data. In [137], the previous federated Bayesian optimization model is improved, where the privacy preservation is reinforced by DP, and a novel distributed exploration technique accelerates the exploration phase of the algorithm. In [19], a digital twin-assisted FA scheme is proposed for Bayesian optimization on distribution estimation. It utilizes an FA version of MCMC combined with a rejection–acceptance sampling technique and a delayed rejection technique. In [85], the observation values of the clients in the federated Bayesian optimization system are protected with LDP satisfaction based on the Laplace mechanism. In [138], a trend prediction problem is formulated by the Bayesian optimization problem and is resolved in an FA model. In that model, observations in each client are summarized by a distribution described by a feature vector. Trend prediction is conducted by analyzing the distributions over a time series. To aggregate the distributions, secure aggregation is performed on the feature vectors based on secret sharing. In [139], a solution for federated Bayesian network structure learning is proposed. The problem is formulated as continuous optimization and is resolved by the FA version of the ADMM algorithm so that only model parameters are exchanged.
IX-C Matrix factorization
TABLE IX: Summery of federated matrix factorization solutions.
Reference Type Separation Privacy Extra features
Dolui et al. [140] Embedding Horizontal DP & Data anonymization Application on a recommender system
FedMF [56] Embedding Horizontal Homomorphic encryption Gradient-based mathod
F2MF [141] Embedding Horizontal DP Fairness concern
Li et al. [57] Embedding Horizontal & Multiple rows & Discrete entries DP & Secure aggregation & Gradient clipping Abstracted computation paradigm for three data separations
Gao et al. [142] Embedding Horizontal & Discrete entries Null Analysis on privacy attack models; Auxiliary data setting
Grammenos et al. [143] SVD Multiple rows (streaming data) DP Time-independent & asynchronous
FedSVD [58] SVD Multiple rows Secure aggregation Lossless SVD results
FedPower[144] SVD Horizontal DP Orthogonal Procrustes transformation; Power method
Hartebrodt et al. [145] SVD Multiple rows Null No eigenvector uploading; Preventing information leakage via smart initialization and data approximation
(a) Horizontal
(b) Multiple rows
(c) Discrete entries
Figure 8: Illustration of different data separation settings in federated matrix factorization
Matrix decomposition, or matrix factorization, refers to the procedure of decomposing one data matrix $M$ into multiple output matrics so that the product of the output matrics is exactly (or approximates) the data matrix. The matrix decomposition problem has many variations, categorized by the required characteristics of the output matrics.
Matrix embedding factorization [^4] Many recommender system researchers directly use the term “matrix factorization” to denote matrix embedding factorization considers a data matrix $M$ with shape $m\times n$ ($m$ rows and $n$ columns), and derives two output matrices $A$ with shape $m\times k$ and $B$ with shape $k\times n$, i.e.,
$M[m\times n]=A[m\times k]B[k\times n].$ (13)
Matrix embedding factorization has many important applications in recommender systems and collaborative filtering, as it directly models the recommendation problems of production environments. Consider a recommender system setting with $m$ consumers (users) and $n$ products for advertisement (items), and the preference of user $i$ towards item $j$ is recorded by the $(i,j)$ entry of the data matrix $M$. The matrix embedding factorization algorithms output the user embedding matrix $A$, and the item embedding matrix $B$, where each row of $A$ is the $k$-length embedding of one user, and each column of $B$ is the $k$-length embedding of one item. With matrix embedding factorization, each user and item can be summarized by a fixed-length vector, representing their characteristics in the user-item interaction. As $k$ is usually set to be much smaller than $m$ and $n$, the equation $M=AB$ can not be guaranteed to have a solution. Therefore, the algorithm designers instead output $A$ and $B$ that minimize the square error in deriving user-item preferences, i.e.,
$\min_{A,B}\sum_{M(i,j)\in M}(M(i,j)-A(i,:)B(:,j))^{2},$ (14)
or minimize the normalized square error, i.e.,
$\min_{A,B}\sum_{M(i,j)\in M}(M(i,j)-A(i,:)B(:,j))^{2}+\lambda(||A||^{2}+||B||^{2}).$ (15)
The aforementioned practical scenario setting of matrix embedding factorization naturally yields the need to solve it in a federated way. Existing federated matrix factorization solutions can be roughly categorized into three categories, based on how the data matrix $M$ is separated among clients: horizontal setting, multiple rows setting, and discrete settings. In the horizontal setting, each client exactly possesses one row of the $M$. In the multiple rows setting, each client possesses multiple rows of $M$ [^5] The term “row” in the two settings can be replaced by “column”, without loss in correctness. In the discrete entries setting, one client cannot obtain a complete row. Information of any complete row is separated by multiple clients.
Horizontal federated matrix factorization originates from the setting that the FA system treats the users as the clients which exactly hold one row of the data matrix $M$ [140, 56, 141]. [^6] This is equivalent to the case that each client holds one column of $M$. In [140], the authors propose a federated matrix factorization algorithm that considers the common case that each client exactly holds all ratings of one user (one row of $M$). Their proposed algorithm lets clients compute the gradients on $A$ and $B$, and transmit the gradients instead of raw data. In [56], another federated matrix factorization solution FedMF is proposed considering a similar problem setting. It also applies gradient transmission, and utilizes homomorphic encryption to prevent the privacy risk caused by exposing the gradient to the server. In [141], the user group fairness – fairness of recommendation accuracy among different user groups – is considered in the federated matrix factorization problem. The authors design a DP scheme to communicate the matrics needed for computing the group fairness.
Other data separation settings of federated matrix factorization are also investigated. In [57], in addition to the aforementioned horizontal setting, the authors consider the setting where each client possesses ratings of a subset of users (multiple rows setting), and the setting where each client possesses all users’ rating on a subset of items (discrete entries setting). The authors then propose FMF as a unified framework to resolve all these data separation settings, with the embedding clipping technique to ensure DP. In [142], the authors consider the setting where each client possesses partial item ratings on a subset of users (discrete entries setting), and the setting where some clients possess auxiliary data of user-item interaction. The authors then discuss the privacy threats and available privacy-preserving approaches in these settings.
Singular value decomposition (SVD) is another important variation of matrix factorization. For a data matrix $M$ with shape $m\times n$, SVD decomposes $M$ into the product of three output matrics $U,V$ and $\Sigma$, so that
$M=U\Sigma V,$ (16)
where $U$ with shape $m\times m$ is a complex unitary matrix, $\Sigma$ with shape $m\times n$ is a rectangular diagonal matrix with non-negative values (singular values of $M$) on the diagonal, and $V$ with shape $n\times n$ matrix with its conjugate transpose $V^{*}$ to be a complex unitary matrix.
As a fundamental matrix factorization technique, SVD has many important applications in data analytics, such as recommender systems. Here we make a similar example to that of matrix embedding decomposition. Consider that $M$ stores user-item rating data with $m$ users and $n$ items. The SVD results are also able to derive vertex-type representations of the users and items. Essentially, each row of $U$ is the embedding of one user, and each row of $V^{*}$ is the embedding of one item.
Federated SVD solutions are then proposed to tackle the scenarios when the data in $M$ is held by distributed clients. These solutions output the SVD results while preserving the privacy of client-held input data. In [143], the authors propose a federated principle component analysis algorithm, whose outputs are exactly approximations of the SVD results. The authors consider the scenario where each client possesses some rows of the data matrix $M$. The local data of each client is generated in a streaming fashion, which leads to the challenge of limited local storage. The proposed algorithm includes tailor-designed local computation and aggregation schemes and satisfies DP as a formal privacy guarantee. In [58], FedSVD is proposed as another federated SVD solution. It relies on a trusted authority that distributes masks to the clients. The clients add masks to their data and upload the masked data to the server with secure aggregation. The masks are delicately designed for SVD so that users can easily unmask the computation results received from the server. With FedSVD, SVD results can be derived accurately without privacy preservation.
In [144], FedPower is proposed as a federated eigenspace estimation (part of SVD procedure) algorithm. It designs a distributed version of the power method, a classical eigenspace estimation scheme, and leverages the orthogonal Procrustes transformation to improve the aggregation phase of the federated power method, to resolve the orthogonal ambiguity. Gaussian noise is added to the possessed matrix in each iteration to enforce Rényi DP. In [145], a federated SVD solution is proposed for genome-wide association studies. Their solution is based on the Gram-Schmidt algorithm. The authors emphasize the advantages of genome-wide association studies, as the computation overhead is independent of the sample size, only depending on the size of features.
(a) Server-client architecture with global model
(b) Server-client architecture with personalized model
(c) Fully decentralized architecture
Figure 9: Illustration of three types of federated MAB settings
TABLE X: Summery of federated MAB solutions.
Reference Architecture Privacy Extra features
FMAB [59] Server-clinet Null Global model; Reward heterogeneity
PF-MAB [60] Server-clinet Null Personalized model; Reward heterogeneity
Li et al. [61] Both DP Global model; Privacy-regret trade-off
Zhu et al. [62] Decentralized DP Personalized model
FedUCB [63] Both Federated DP Global model; Contextual bandit
IX-D Multi-armed bandits
The multi-armed bandits (MAB) model considers a system with $k$ arms, each with a fixed distribution of rewards, and an agent successively pulling the arms. Each time the agent selects an arm and pulls it, a stochastic reward determined by its corresponding reward distribution is returned to the agent. The agent aims at maximizing the reward it received, or formally speaking, minimizing the cumulative regret $\rho$:
$\rho=T\mu^{*}-\sum_{t=1}^{T}r_{t}.$ (17)
The aforementioned definition $\rho$ indicates the cumulative regret after $T$ pulls, where $\mu^{*}$ indicates the maximal reward mean among all arms, and $r_{t}$ indicates the received reward in the $t$-th pull. An MAB algorithm performs decision-making in selecting the arm to be pulled in each round. The algorithm handles the famous exploration-exploitation tradeoff: whether the agent should select an underexplored arm with the potential to have a high reward mean, or should exploit an arm that has a relatively high reward mean based on previous trials.
As the effectiveness of an MAB algorithm highly relies on gathering and analyzing previous trial results, the federated MAB model emerges to enable multiple clients (MAB agents) to collaboratively improve their own MAB performances. The federated MAB model assumes that multiple clients are having their MAB models with a shared set of arms. By sharing the trial results, the MAB training can be greatly improved.
There exist at least three kinds of federated MAB setting, as demonstrated in Fig. 9: 1) Server-client architecture with global model, where a single server model is optimized by aggregating the trial results from all clients; 2) Server-client architecture with personalized model, where each client processes its own MAB model, and optimizes its local model by leveraging both trial results from itself and trail results of other clients received from the server; 3) Decentralized architecutre, where personalized models are also trained by the clients, but no centralized server exists to aggregate all trial results. Instead, each client can only receive trial results from a portion of clients via client-to-client communication.
In [59], a federated MAB solution is proposed. It considers the scenarios that the server would like to learn the global bandit model, while the MAB model of the clients is diverged from the global model regarding the arm reward means. To tackle the challenge of diverged MAB models, in addition to the basic federation scheme which transmits the trial results between server and clients, the authors propose a novel confidence bound-based scheme, named Fed2-UCB, to handle the uncertainty of sampling clients with the heterogeneous arm reward means. In [60], another federated MAB solution is proposed to handle a different federated MAB setting. The new solution also handles the problem of heterogeneous reward means. In addition to the global bandit model, in [60], each client would like to train its personalized MAB model that fits its local environment. By carefully leveraging the knowledge from other clients, their proposed federated MAB algorithm, PF-MAB, achieves good regret bounds for each client, despite that the information from other clients might be misleading. In [61], the authors consider enhancing privacy preservation in transmitting the trial information. They apply DP to the uploads and modify the UCB algorithm of the MAB model to handle the uncertainty introduced by DP. Their system is deployed in both the server-client architecture and the fully decentralized architecture, where clients independently communicate with other clients and share information about the trial results. In [62], the authors further investigate the fully decentralized federated MAB models. The proposed Gossip-UCB algorithm tackles the coupling effect of client gossiping, considers the effect of biased client rewards, enforces DP, and provides a theoretical guarantee on regrets. In [63], the authors investigate the federation of contextual bandits, a variation of MAB with parameterized trials and arms. In their proposed scheme, the Gram matrix of actions and reward vector, which are the information about the regression of rewards, are transmitted. The proposed FedUCB algorithm, based on the LinUCB algorithm for centralized contextual bandits, handles both the server clients and decentralized architectures and enforces DP for better privacy preservation.
IX-E Unsupervised representation learning
Unsupervised representation learning refers to the procedure of deriving expressive representations (features) for data in various structures, where the data are usually unlabelled. It converts the raw data with complex structure and high dimensionality into vectors with limited dimensions, which are feasible for many downstream tasks. Unsupervised representation learning is also applied in federated scenarios, where the data stored in multiple clients are given their features in a common scope. The federated representation learning problem is often resolved in the FA paradigm, where clients upload insights about their local data and features, and the server aggregates the insights to derive a comprehensive view of the data distribution and feature space. Many representation learning algorithms utilize deep learning, and their federated versions utilize FL. However, FA is still needed in these algorithms as the complex information exchange cannot be accomplished by simply averaging the parameters of neural networks.
In [146], a federated representation learning algorithm that utilizes both FA and FL is proposed. It realizes unsupervised representation learning via contrastive learning. It trains a feature learning model via FL and utilizes FA to exchange local features with other clients. With the knowledge of global feature distribution, contrastive learning can be properly performed to derive distinguishable features, even if the data within each client is highly skewed. In [147], an unsupervised federated representation learning method FedCA is proposed. In addition to the contrastive learning model trained by FL, FedCA proposes a dictionary module trained by FA to align the learned representations among clients with non-IID data. The dictionary model is first derived locally by adding the normalized representations of local data samples and then aggregated by the server to derive the global dictionary. FedWalk [148] is an FA-based solution for graph embedding analytics. It derives expressive representations for the vertexes in a large graph, which has a complex and uncertain data structure. FedWalk considers the vertex-separated federated setting, where each client represents one vertex in the graph, and only possesses neighborhood information of the vertex. It performs the FA version of a random walk, a classical graph embedding learning algorithm.
X Assisting federated learning
Figure 10: A typical workflow of FL assisted by FA.
FL has gained success in various data-incentive tasks where a neural network is trained for prediction/description applications. Many works are proposed to improve the FL performance regarding model accuracy, convergence rate, scalability, etc. In these works, the data of the clients are usually analyzed to pursue optimization of the FL performance. However, as is restricted by the tenet of FL, the raw data should not be transmitted to the server while conducting such analysis. Therefore, in these works, the raw data of the clients are usually transformed into insensitive insights which help the system to perform decision-making in optimizing FL. Although not explicitly claimed, such an approach exactly follows the FA paradigm, and these solutions can be regarded as an FL instance optimized by another FA algorithm.
We categorize the solutions of FA-assisted FL based on when FA takes effect, as illustrated in Fig. 10. We summarize the whole procedure into three phases: the preprocessing phase, where the clients are registered and evaluated by the server; the training phase, where the FL model is transmitted between server and clients, trained by the clients locally, and aggregated by the server, for multiple rounds; the postprocessing phase, where the performance of the derived FL model is then evaluated. Strictly speaking, FL only includes the training phase, and the assisting FA schemes can be executed in the preprocessing phase, in the postprocessing phase, and sometimes in the training phase parallel with FL. Fig. 10 illustrates the involvement of FL and FA in the whole procedure with colors.
In the following parts, we will introduce some classes of FA schemes to assist FL. In Section X-A, we investigate FA schemes that help FL to select proper participating clients, which executes on the preprocessing phase and client selection step of the training phase. In Section X-B, we investigate FA schemes that help FL model personalization via conducting clustering on the clients, which executes on the preprocessing phase and model aggregation step of the training phase. In Section X-C, we investigate FA schemes to evaluate the performance of the FL model, which executes in the postprocessing phase.
X-A Preprocessing: Client selection in FL
In FL, only part of the clients will participate in each round of FL, due to the limits of device availability and system capacity. Many pieces of research show that clients provide different benefits/contributions to the model training determined by the data size and quality of the clients. Therefore, wisely selecting the participating clients in each round then has the potential to obtain a better FL performance, compared to the default setting of random selection. Selecting the proper participating clients requires analyzing the characteristics of the client data, which naturally requests the power of FA.
FedACS [8] is an FA solution to measure the data skewness (severity of data heterogeneity) of the clients. Since data heterogeneity is a significant source of performance degrading of FL tasks, measuring data skewness is proven to be beneficial in assisting client selection for FL. In FedACS, the FL gradients of clients are utilized as FA insights, so that the computation and infrastructure of FL can be reused. The authors prove that the clients with local gradients close to the average gradient tend to have lower data skewness, and are more beneficial to the FL model. The server then formulates a dueling bandit to perform client selection, where clients with lower data skewness are more likely to be selected. In FAVOR [149], the authors formulate a reinforcement learning model running on the server side to perform client selection. In that reinforcement learning model, the states are defined to be a combination of the global model gradient and local gradients of all clients, where the latter is transmitted from the clients and serves as the FA insight. The actions are defined as all possible sets of participating clients in the incoming round. The rewards are defined by the validation accuracy of the model after the training. Such reinforcement learning formulation forces the reinforcement learning model to choose the participating clients that achieve higher validation accuracy. In Oort [150], the authors optimize the training and evaluation of FL models by wisely selecting participating clients. The optimization of training is based on an observation of the authors, that a client inducing higher training loss is more beneficial for the FL training. Therefore, in Oort, the clients upload their training loss (which is a scaler) to the server in each round they participate in, which works as the insight of FA. The server then utilizes a heuristic multi-armed bandit to perform client selection, where clients obtaining higher training losses in previous rounds will be more likely to be selected. In [151], the authors formally prove that selecting clients with higher training loss can accelerate the FL model convergence. Based on the theoretical result, the authors propose a power-of-choice client selection strategy, where a global model is sent to all available clients, the clients upload the training loss as the FA insight, and the server finally selects a set of high-loss clients as the participants in the following several rounds.
X-B Preprocessing: Personalize FL models via client clustering
The vanilla setting of FL considers an identical global model shared by the server and clients, which is applied in downstream tasks on all data. However, in practice, the client data are usually non-IID distributed, and the data held by different clients has quite different characteristics (distributions). As a result, it becomes difficult to let one model achieve good generalization on data from all clients, i.e., the optimal model parameters for global data might be quite divergent from the optimal model parameters for data in particular clients. Personalized FL resolves such an issue by breaking the limit of one identical global model. In personalized FL, multiple models are trained by data from different clients and are applied to the data of those clients. With personalized FL, each client can obtain a personalized model tailored to its data characteristics, which achieves better performance when performing inference on its data [152].
Among various approaches to personalized FL, clustering-based methods are widely selected for their high reasonability and strong performance. In clustering-based methods, clients are firstly clustered, where those with similar data characteristics are placed in the same cluster. Then, FL is fine-tuned based on the data within one cluster. With such approaches, the personalized models will be trained on sufficient data stored in a number of clients, and tailored for particular data characteristics. The clustering procedure requests the sharing of information about the client data. To preserve privacy, existing clustering-based personalized FL algorithms exactly follow the FA paradigm in clustering the clients. The clustering problems can be regarded as a special kind of clustering problem in general data analytics, and those client clustering algorithms can also be regarded as a special kind of FA-based clustering solutions.
In [9], the authors propose a pioneering clustering personalized FL solution named CFL. In CFL, the cluster assignment of clients is determined by a hierarchical clustering structure: the data distribution similarity of pairs of clients is calculated by the cosine similarity of local gradients; a bipartition of clients is derived by analyzing the similarity matrix; more clusters are derived by hierarchically executing the bipartition algorithm as long as it could improve the overall performance. In [153], the authors propose another pioneering clustering personalized FL solution. It trains a global model with vanilla FL and then uses gradients from all the clients to perform clustering. The clustering is based on the cosine similarity between clients using an agglomerative clustering algorithm. In [154], the authors propose a clustering personalized FL approach by extracting one layer of the gradient uploaded by the clients and clustering clients by calculating the cosine distance of the layer parameters. The authors also propose a decomposition and consolidation scheme to remove some extreme clusters and reassign the clients to other clusters. In [155], the authors propose an Iterative Federated Clustering Algorithm (IFCA) to cluster the clients for personalized FL. In IFCA, the server first formulates one model for each of the clusters and distributes all the models to the participating clients. The clients calculate their local loss of all the models, and one client will be assigned to the cluster where the lowest loss is obtained on the corresponding model. In [156], the authors propose COMET to perform clustered co-distillation in FL. In COMET, there exists a public set of unlabeled data samples. The clients train their personalized model with its local data and then infer the public data to derive the soft decisions. The clients upload their soft decisions on the public data to the server, and the server performs $k$-means on that so that clients with similar soft decisions on the public data will be placed in the same cluster. After that, knowledge sharing is performed in model training. In detail, the clients then train their local model with their local data, while an extra regularization term is applied. The regularization term forces the model to obtain similar soft decisions to other clients in the same cluster on the public data. In [157], the authors consider the problem that the clients may find it hard to cluster into several distinct data distributions. They consider soft clustering, where the distribution of data in each client is considered as a mixture of several clusters. FedSoft is proposed to perform soft clustering by requesting the clients to train with a proximal loss, regulating the training to get close to a mixture of several cluster models. FedSoft system employs an FA subroutine to estimate the weight of each client with all the clustered. Each data sample of one client is evaluated by all the cluster model, and is assigned to the cluster with the smallest loss. The weight of the client is then calculated as the portion of data samples belonging to all the clusters.
X-C Postprocessing: Model evaluation
While the mainstream of FL studies focuses on model training, i.e., deriving a model that is expected to have high performance based on the federated data. Meanwhile, as the evaluation of the trained model also relies on the vast data, leveraging the federated data then becomes an effective approach. In the blog [7] where the term FA is coined, the authors claim that the idea of FA is exactly motivated by the need to evaluate the trained FL model. In Oort [150], the authors propose a system of federated evaluation. Oort automatically adjusts the number of participating clients, so that the accuracy of performance evaluation is guaranteed and the client effort is minimized. Such adjustment is achieved by the confidence bounds derived by Hoeffding’s inequality.
XI Wireless Network Applications
Wireless networks have become a key cornerstone of modern communication systems. Meanwhile, data analytics (machine learning) approaches are widely applied to optimize the performance of wireless network systems. These technologies have been utilized to handle critical tasks such as network slicing, caching management, anomaly detection, and semantic communication. Federated computation, including both FL and FA, is also applied by the researchers to obtain the benefits of privacy preservation, each computing utilization, and communication efficiency. In this section, we focus on the FA solutions applied for the critical optimizations on wireless network systems.
Some FA solutions directly enable non-neural network data analytics powers for wireless network challenges. They design federated versions of data analytics algorithms, such as clustering, to optimize wireless network systems. In [158], an FA solution is proposed for anomaly detection in cellular network antenna tilt. It transforms the data regarding antenna electrical tilt into signature vectors, and then performs federated clustering on these signature vectors, where the local clustering centroids serve as the FA insights. Then, anomaly items are then filtered out by measuring the distance between signature vectors and global clustering centroids. In [159], the problem of energy consumption minimization in mobile edge computing task offloading is considered. To handle the time-varying user task, a federated version of support vector machine (SVM) is proposed. In the federated SVM scheme, the covariance matrix between users is transmitted from server to clients (base stations), and the SVM weights are uploaded from clients to server, and are then aggregated by the server. Eventually, the global SVM can derive the optimal solution of task offloading, with optimized communication and computation overheads. In [160], an FA scheme is utilized in mobile crowdsensing scenarios to discover malicious sensing tasks. In such a scheme, a machine learning module is equipped by each client (detection device). The clients predict whether a task is malicious, and uploads the prediction results to the server as the FA insight. The server aggregates the insights, and derives the updated reputation score of the task proposers to the clients, which helps optimize the prediction results in the future. The global prediction results are also derived by aggregating the client predictions, with extra risk-aware computations.
On the other hand, other FA solutions optimize the wireless network systems based on an FL backbone, i.e., , a neural network is trained collaboratively to directly optimize the system. The FA algorithm then proposes assistance on the FL scheme. Such methodology has been applied in numerous FA research, as is introduced in Section X. In [161], FL is employed to build up a semantic communication system, while FA is applied to assist it via generating the aggregation indicators. The aggregation indicator is generated by the FA local computation scheme, reflecting the suitability between the current channel condition and the local FL model. When the aggregation indicator is received by the server along with the FL model, it is transformed into the weight in weighted FL model aggregation. In [162], the FA-assisted FL scheme is deployed to control the UAV deployment in visible light communications scenarios. It needs to resolve the complex optimization problem jointly considering “UAV deployment, user association, power efficiency, and predictions of the illumination distribution”. To predict the illumination distribution, a convolutional autoencoder is trained by FL methodology. To help the FL training, the FA methodology is also applied. The FA scheme uploads the convolution kernels and bias to the server. By aggregating the convolution kernels, gap matrices can be calculated, which are essential for the evaluation and optimization of the UAV deployment plan. In [163], an FL system is designed to conduct intrusion detection. To tackle the challenge of privacy risk of transmitting gradient, non-IID data, and high communication overhead, the proposed scheme exactly replaces the conventional FL insight derivation and aggregation schemes, by those in FA methodology. Such schemes, utilizing the idea of knowledge distillation, are also termed federated distillation from the FA perspective. In the proposed scheme, the locally generated labels on shared unlabelled data are sent to the server, and then aggregated to form the global hard labels. Then, the server can use the shared data with global labels to train the global model. In [164], an FL model is trained to perform content popularity prediction in Fog-RAN systems. The clustered FL scheme is further deployed based on the federated clustering solution. In the federated clustering scheme, the local features of the clients are computed and uploaded. The server computes the similarity between clients, and splits up new clusters iteratively. Eventually, the final clustering results can be derived, and the personalization of the host FL scheme is improved.
XII Open issues and future directions
XII-A Applications to more complex data scenarios
The fundamental difference between FA and FL lies in the nature of their supporting tasks. The scope of data scenarios studied by existing FA remains small compared with the diversity of data science problems and models in the wild. Designing FA mechanisms to support non-trivial data types, including graph data [165], streaming data[166], key-value data[167], multidimensional data [168], time series data, and important networked applications, including Internet telemetry[169], smart home [170], healthcare [11], web 3.0 and other privacy-critical senarios[171] all present interesting and open challenges.
XII-B Privacy preservation at scale
As FA systems grow in scale with more data or users involved, ensuring privacy becomes more challenging. This challenge is exacerbated when analysts seek high-accuracy analytics. Existing privacy-preserving techniques exhibit distinct characteristics concerning privacy assurance, data utility, and scalability. For example, differential privacy has long been criticized for offering a poor accuracy-privacy tradeoff, prohibiting its applications in domains that request accurate population-level profiling or analytics. Other cryptographic techniques, such as multi-party computation, encounter difficulties in scaling to practical sizes for mobile applications. Consequently, FA mechanisms have to be carefully designed to strike a delicate balance among privacy, utility, and scalability.
XII-C A unified FA framework
FL operates under a unified framework to embed similar computation procedures and insight structures when training various types of neural networks. Such a unified framework plays a vital role in boosting the wide studies and applications of FL. On the other hand, FA mechanisms designed so far are still highly task-specific. While it seems intuitive that different data analytics tasks are naturally different regarding their computation procedures, data structures, etc, there already exists some efforts to provide a unified framework for a particular class of data science problem[16]. Attracted by the potential benefits of a unified framework, it remains a grand open problem to design a unified framework for broader classes of data analytics tasks or even universal data analytics tasks. Specifically, motivated by this, some open questions and interesting directions could be (i) What are the proper architecture, and interface design for a unified FA framework? (ii) If we still adopt the current local insight upload and global aggregation scheme, how to design a general aggregator or general insight form that capable of incorporating diverse requirements? (iii) Witnessing the advancement in generative models, is it possible to use generative models at the server side, essentially, as a general aggregator, to further reduce the reliance on locally uploaded insights so that the communication cost can be reduced further and versatile privacy-preserving tasks can be supported.
XII-D System-efficient FA
Studies in the field of FA have primarily concentrated on algorithm design to enable the execution of various data analytics tasks in a federated manner. In contrast to the well-established problem hierarchy in FL, it is evident that a substantial portion of the FA domain remains unexplored, particularly from the systems’ perspective. As a federated system, FA also faces practical challenges, such as heterogeneous computing power across devices, dynamic system sizes with devices joining and leaving, incentives and pricing issues, robust FA under adversaries, etc. FA further introduces new interpretations of system measures, such as fairness, and novel system challenges, such as significant computing costs for certain privacy-preserving measures, and a limited privacy budget on the client side. As FA continues to evolve, addressing these system-level challenges will be crucial for its successful implementation and widespread adoption.
XII-E Wireless communication for FA
Over-the-air computation (AirComp) leverages the superposition property of wireless channels to enable simultaneous transmission and aggregation of signals from multiple devices [68]. It has been applied to FL by aggregating model updates from numerous devices directly over the air. It not only enhances communication efficiency but also maintains the privacy of clients’ local data, addressing key bottlenecks in traditional FL approaches, such as limited bandwidth and latency issues. How FA could benefit from physical layer techniques also presents interesting and open challenges.
XIII Conclusion
Due to the exponential growth of edge data and the growing awareness of data privacy, privacy-preserving distributed data processing has attracted wide interests from both academia and industry. Federated analytics is an emerging collaborative data processing framework for descriptive data science tasks without centralizing the raw data. It brings significant benefits in privacy protection, communication reduction, and task coverage. Although FA has been widely studied in industry and academia, a systematic review of the existing efforts in FA has not been conducted yet. This survey fills the gap by first comprehensively reviewing the key concepts in FA and its relationship with similar techniques. It then presents a detailed taxonomy to categorize FA studies from both applications and system characteristics. Key challenges and enabling techniques are introduced with a specific focus on privatization, analytics, and optimization. A wide spectrum of FA tasks are then reviewed demonstrating the generality of FA applications. Finally, we discuss the open issues and future directions in FA, from the perspective of application, privacy protection, framework, system optimization, and cross-layer design. This survey summarizes the existing efforts in FA and the huge intersection among FA and data science, privacy, distributed computing, wireless communications, and networking systems. Overall, FA research is widely interdisciplinary, approachable, and contains many critical problems that would benefit greatly from the expertise of all related areas.